### v1

# Universal Approximation Theory: The basic theory for large language models

Abstract
Language models have emerged as a critical area of focus in artificial intelligence, particularly with the introduction of groundbreaking innovations like ChatGPT. Large-scale Transformer networks have quickly become the leading approach for advancing natural language processing algorithms. Built on the Transformer architecture, these models enable interactions that closely mimic human communication and, equipped with extensive knowledge, can even assist in guiding human tasks. Despite their impressive capabilities and growing complexity, a key question remains – the theoretical foundations of large language models (LLMs). What is it about the Transformer architecture that makes it so effective for powering intelligent language applications, such as translation and coding? What underlies LLMs’ ability for In-Context Learning (ICL)? How does the LoRA scheme enhance the fine-tuning of LLMs? And what supports the practicality of pruning LLMs? To address these critical questions and explore the technological strategies within LLMs, we leverage the Universal Approximation Theory (UAT) to offer a theoretical backdrop, shedding light on the mechanisms that underpin these advancements.
Universal Approximation Theory: The basic theory for large language models
Wei Wang1, Qing Li1, 1The Hong Kong Polytechnic University, weiuat.wang@connect.polyu.hk, qing-prof.li@polyu.edu.hk
1 Introduction
In recent years, the rapid emergence of LLMs in the field of artificial intelligence has undoubtedly become one of the most notable advancements within the domain. The core allure of these models stems from their extraordinary capabilities in language processing. Language, as a unique crystallization of human intelligence, serves not only as the external reflection of thought but also as the bridge for communication, the cornerstone for the dissemination of knowledge, and the continuation of civilization, profoundly shaping the identity of humans as a unique species. Thus, endowing machines with the ability to understand and generate language marks a significant leap towards the realization of true artificial intelligence. The emergence of models such as the ChatGPT Radford and Narasimhan (2018); Brown et al. (2020); Achiam et al. (2023), the Llama Touvron et al. (2023), and the PaLM Chowdhery et al. (2023) vividly demonstrates this point.
A distinctive feature of LLMs is their immense parameter size Achiam et al. (2023); Touvron et al. (2023); Chowdhery et al. (2023); Chen et al. (2021); Zeng et al. (2022), often amounting to hundreds of billions or even trillions (for instance, GPT-3’s Brown et al. (2020) 175 billion parameters and PaLM’s Chowdhery et al. (2023) 540 billion parameters). This vast parameter scale lays the foundation for their exceptional language processing capabilities and enables them to exhibit almost human-like traits, such as ICL  Brown et al. (2020); Dong et al. (2022), instruction following Sanh et al. (2021); Ouyang et al. (2022); Wei et al. (2021), and multi-step reasoning Wei et al. (2022). Notably, these colossal models are predominantly trained by tech giants like Google and Microsoft using large-scale GPU clusters Zhao et al. (2023), sparking a research fervor on how to efficiently fine-tune them with limited GPU resources. The advent of Lora Hu et al. (2021) fine-tuning technology has provided an effective pathway for this, allowing for the fine-tuning of large models under resource constraints without the need to adjust all parameters of the original model comprehensively. Moreover, model pruning techniques Sun et al. (2023); Ma et al. (2023) are crucial for deploying large models in resource-constrained environments, aiming to reduce the model size for operation on smaller devices. Faced with the challenge of processing long texts, such as generating summaries or answering questions based on extensive documents—which traditionally requires substantial computational resources—technologies like LongLora Chen et al. (2023) have been developed to tackle the difficulties of processing long contextual texts, further expanding the application boundaries of LLMs. Today, LLMs possess a diverse range of functionalities, from translation and text summarization to automatic code generation, demonstrating their versatility.
Despite LLMs being on a fast track of development towards higher intelligence and reliability, the theoretical foundation behind them remains largely unexplored, shrouded in clouds of uncertainty. The scientific community is actively seeking to uncover the inner mechanisms behind their powerful capabilities, including analyses of ICL mechanism Xie et al. (2021); Min et al. (2022), among others. Against this backdrop, we establish the UAT Cybenko (2007); Hornik et al. (1989) as the mathematical essence of LLMS, and utilize the UAT lens to clarify pivotal technologies and phenomena within contemporary LLMs, aiming to illuminate this central theoretical conundrum from a fresh perspective. Our contributions are as follows:
  • We prove that the Transformer is the tangible embodiment of UAT.
  • We deliver a rigorous scientific explanation of Transformer-based LLMs through the UAT.
  • We explain the characteristics of LLMs, such as ICL, instruction following, multi-step reasoning, and the technologies applied within LLMs like Lora, pruning, and LLMs’ strong generalization capabilities.
  • We offer insights into the future development of LLMs.
Our article is structured as follows: In Section 2, we begin by introducing the UAT and propose that to demonstrate the Transformer’s adherence to UAT, it is necessary to show that both Linear and Multi-Head Attention (MHA) can be represented in the form of matrix-vector multiplication. In Section 3, we establish that the Transformer belongs to the category of UAT. Our proof strategy unfolds starting with Section 3.1, where we present the idea to express Linear and MHA as matrix-vector products. In Section 3.2, we detail the computational methods required for the proof. Sections 3.3 and 3.4 are dedicated to proving that both Linear and MHA can indeed be represented as matrix-vector operations. Further in Section 4, we leverage UAT to theoretically elucidate some fundamental issues (Generalization 4.1, ICL 4.2) and techniques (Pruning 4.3 and LoRA 4.4) associated with LLMs. Finally, we conclude with a summary of the existing challenges and potential future developments for LLMs in Section 4.5.
2 The Universal Approximation Theory
Up to now, the UAT (Cybenko, 2007) is the most widely recognized fundamental theory in deep learning. However, this theorem is only applicable to the simplest form of neural networks, the multilayer perceptron (MLP) (Cybenko, 2007; Popescu et al., 2009). Due to the increased complexity of Transformer networks, they cannot be mathematically expressed in the same form as UAT. Thus, the theory has not yet been extended to Transformer networks. Our aim in this paper is to unify Transformer networks under the framework of UAT, thereby standardising their mathematical representation. Before unifying their mathematical forms, we provide a brief overview of the theorem, which was initially proposed by Cybenko (2007). This theorem encompasses numerous conclusions and proof details. Although it has been further developed, its fundamental mathematical form remains unchanged. Therefore, this paper explains the theory based on the UAT form presented by Cybenko (2007). Theorem 2 from Cybenko (2007) states that if $\sigma$ is any continuous sigmoidal function, then finite sums of the following form:
$G(\mathbf{x})=\sum_{j=1}^{N}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$ (1)
is dense in $C\left(\mathbf{I}_{n}\right)$. Here, $\mathbf{W}_{j}\in\mathbb{R}^{n}$ and $\alpha_{j},\theta\in\mathbb{R}$ are fixed. For any $f\in C\left(\mathbf{I}_{n}\right)$ and $\varepsilon>0$, there exists a function $G(\mathbf{x})$:
$|G(\mathbf{x})-f(\mathbf{x})|<\varepsilon\quad\text{ for all }\quad\mathbf{x}\in\mathbf{I}_{n}.$ (2)
This implies that, when $N$ is sufficiently large, a neural network can approximate any continuous function on a closed interval. Hornik et al. (1989) further demonstrates that multilayer feedforward networks also conform to the UAT, capable of approximating arbitrary Borel measurable functions. Observing Equation (1), where the function $G(\mathbf{x})$ yields a scalar output in $\mathbb{R}$, the scenario expands naturally when $G(\mathbf{x})$ maps to $\mathbb{R}^{m}$, requiring the approximation in each dimension. It becomes evident that to accommodate this multidimensional output, a simple extension to Equation (1) suffices: the transformation matrix $\mathbf{W}_{j}$ is revised to the space $\mathbb{R}^{n\times m}$, the bias term $\theta_{j}$ is recast as a vector in $\mathbb{R}^{m}$, and $\alpha_{j}$ is reshaped into a diagonal vector. Nevertheless, in these demonstrations, the theorem does not straightforwardly apply to Transformer architectures. However, if we succeed in reformulating both the Linear ($\mathbf{x}|\mathbf{W}$) and MHA ($\mathbf{x}|\mathbf{W}$) components into a unified representation, $\mathbf{W}^{\prime}\mathbf{x}^{\prime}$, where $\mathbf{W}$ and $\mathbf{x}$ represent the parameters and inputs for each component respectively, and $\mathbf{W}^{\prime}$ is derived from $\mathbf{W}$ while $\mathbf{x}^{\prime}$ signifies a column vector (for details, refer to Section 3.1). It is easy to prove that multi-layer Transformer networks are concrete embodiments of UAT.
Consequently, our objective is to represent Linear and MHA architectures as a matrix-vector multiplication similar to Eq. (3). This approach offers two key benefits: Firstly, it allows for a fundamental mathematical investigation of deep learning models. Secondly, it provides a more intuitive and objective basis for comparing differences among various models. On this foundation, we propose to transform Linear and MHA into the matrix-vector format.
3 UAT for Transformer
In the previous section, we have outlined our fundamental goal to unify the Transformer model under the framework of UAT by converting Linear and MHA operations in the Transformer into matrix-vector forms. In this section, we will use the Matrix-Vector method to transform Transformer operations into matrix times vector forms.
3.1 Matrix-Vector Method
Figure 1: The transformation process of the Matrix-Vector Method.
Before delving into transforming Linear and MHA into their matrix-vector format, we introduce the Matrix-Vector Method, which will subsequently be employed to cast both Linear and MHA operations into a unified matrix-vector format. This method constitutes a strategic realignment of input data and corresponding parameters of various transformations within the network, as illustrated in Figure 1. The underlying principle is as follows: both the input ($\mathbf{x}$) and output ($\mathbf{y}$) data are reconfigured uniformly through a transformation $\mathbf{T_{D}}$ into column vectors ($\mathbf{x}^{\prime}$ and $\mathbf{y}^{\prime}$, while parameter tensor ($\mathbf{W}$) is reorganized into matrix form ($\mathbf{W}^{\prime}$). A critical requirement is that identical operations throughout the LLMs network adhere to the same restructuring scheme, thereby eliminating the need for additional transformations on intermediate feature data; they can directly be represented as column vectors. Additionally, by default, matrix variables in the original formulas are represented in bold, such as $\mathbf{x}$, and in the Matrix-Vector form, corresponding variables are denoted with a prime symbol () in the upper right corner, such as $\mathbf{x}^{\prime}$. Elements within matrices are represented by corresponding lowercase letters with subscripts, for example, $x_{i}$. The Matrix-Vector Method can be succinctly encapsulated as follows:
$\mathbf{y}=T(\mathbf{x}|\mathbf{W})\rightarrow\mathbf{y}^{\prime}=\mathbf{W}^{\prime}\mathbf{x}^{\prime}$ (3)
It is obvious that $T_{D}$ and $T_{P}$ are not fixed, we could design various kinds of ways to do those. For convenience, we present a methodology tailored for Linear in Section 3.3 and an approach for MHA in Section 3.4, thereby illustrating the adaptability and application of the Matrix-Vector Method across different components of the Transformer architecture.
3.2 Diamond Matrix Multiplication
While our goal is to transform Linear and MHA modules into a matrix-vector format, this process is complex. In order to describe this transformation process clearly, we propose a new computational method called the Diamond Multiplication Method, denoted by the symbol $\diamond$.
$\left(\begin{array}[]{cccc}w_{1,1}&w_{1,2}&\cdots&w_{1,n}\\ w_{2,1}&w_{2,2}&&w_{2,n}\\ \vdots&\vdots&\vdots&\vdots\\ w_{m,1}&w_{m,2}&&w_{m,n}\end{array}\right)\diamond\left(\begin{array}[]{c}x_{1}\\ x_{2}\\ \vdots\\ x_{m}\end{array}\right)$ (4)
$=\left(\begin{array}[]{c}w_{1}x_{1}+w_{31}*x_{2}+\cdots w_{m,1}x_{m}\\ w_{1,2}x_{1}+w_{2,2}x_{2}+\cdots w_{m,2}x_{m}\\ \vdots\\ w_{1,n}x_{1}+w_{2,n}x_{2}+\cdots w_{mn}x_{m}\end{array}\right)$
Let $\mathbf{W}\in\mathbb{R}^{(m,n)}$ and $\mathbf{x}\in\mathbb{R}^{(m,1)}$, the Diamond Matrix Multiplication is defined as $\mathbf{y}=\mathbf{W}\diamond\mathbf{x}$, where $\mathbf{y}\in\mathbb{R}^{(m,1)}$. The computation procedure involves element-wise multiplication and summation of corresponding elements from left to right across the columns of matrix $\mathbf{W}$ with vector $\mathbf{x}$. The sum of the multiplication of the $i$th column of $\mathbf{W}$ with $\mathbf{x}$ is assigned as the $i$th element of $\mathbf{y}$. A detailed calculation process is provided in Eq. ((4)). The computation process for $\mathbf{y}=\mathbf{x}\diamond\mathbf{W}$ is identical to that of $\mathbf{W}\diamond\mathbf{x}$, implying that $\mathbf{W}\diamond\mathbf{x}=\mathbf{x}\diamond\mathbf{W}$. However, it should be noted that the Diamond Matrix Multiplication does not possess the properties of associativity and distributivity.
Furthermore, based on its computation method, it can be inferred that the Diamond Matrix Multiplication is related to conventional matrix multiplication, specifically: $\mathbf{W}\diamond\mathbf{x}=\mathbf{W}^{T}\mathbf{x}$. Additionally, when $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are square matrices, we can derive the following relationships: $\mathbf{W}_{1}\diamond[\mathbf{x}\diamond\mathbf{W}_{2}]=\mathbf{W}_{2}^{T}\diamond\mathbf{W}_{1}\diamond\mathbf{x}$ and $[\mathbf{W}_{1}\diamond\mathbf{x}]\diamond\mathbf{W}_{2}=\mathbf{W}_{1}^{T}\diamond\mathbf{W}_{2}\diamond\mathbf{x}$. Detailed derivations can be found in Appendix A.
3.3 Matrix-Vector Method for Linear
Figure 2: The process of converting a linear transformation into its corresponding matrix-vector representation. a: Depicts the general form of a linear transformation. b: Presents a straightforward example of a linear transformation. c: Demonstrates the transformation of the linear operation from b into the matrix-vector format.
In this section, we present a way to transform Linear operation into matrix-vector format. Figure 2 illustrates this process: Figure 2.a shows the linear transformation of multi-channel input: $\mathbf{W}\mathbf{x}=\mathbf{y}$. Figure 2.b provides a specific example of Figure 2.a, Figure 2.c converts the linear transformation in Figure 2.b into the corresponding matrix-vector representation: $\mathbf{W}^{\prime}\diamond\mathbf{x}^{\prime}=\mathbf{y}^{\prime}$. Thus, the linear transformation can be represented in matrix-vector form as follows:
$\mathbf{x}_{i+1}=\mathbf{W}_{i}\mathbf{x}^{i}\rightarrow\mathbf{x}_{i+1}^{\prime}$ $=\mathbf{W}_{i}^{\prime}\diamond\mathbf{x}_{i}^{\prime}=(\mathbf{W}_{i}^{\prime})^{T}\mathbf{x}_{i}^{\prime}$ (5)
Here, $\mathbf{x}_{i}\in\mathbb{R}^{(N,M)}$ and $\mathbf{x}_{i+1}\in\mathbb{R}^{(N,M)}$ represent the input and output of layer $i$, respectively, while $\mathbf{W}_{i}\in\mathbb{R}^{(N,N)}$ represents the parameters of layer $i$. $\mathbf{x}_{i}^{\prime}$, $\mathbf{x}_{i+1}^{\prime}$, and $\mathbf{W}_{i}^{\prime}$ are generated based on $\mathbf{x}^{i}$, $\mathbf{x}^{i+1}$, and $\mathbf{W}_{i}$ using the Matrix-Vector Method. The derivation of the general form can be found in Appendix B.1.
Figure 3: The Transformer process.
Figure 4: The process of transforming $\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}}_{8})$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{WH}^{\prime}\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$.
Figure 5: This diagram illustrates the process of transforming $\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}})\mathbf{W}_{O}$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{WHO}^{\prime}\diamond\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, different colors in the diagram correspond to different variables.
3.4 Matrix-Vector Method for MHA
We now employ the Matrix-Vector Method to elucidate the inner workings of the MHA. The mechanism is defined by the following equation:
$\mathring{\mathbf{H}}=\operatorname{MultiHead}(\mathbf{Q},\mathbf{K},\mathbf{V})$ (6)
$=$ $\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}$
$\operatorname{Attention}(\mathbf{x}_{i}\mathbf{W}_{Qi}=\mathbf{Q}_{i},\mathbf{x}_{i}\mathbf{W}_{Ki}=\mathbf{K}_{i},$ (7)
$\mathbf{x}_{i}\mathbf{W}_{Vi}=\mathbf{V}_{i})=\operatorname{softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{T}}{\sqrt{M}}\right)\mathbf{V}_{i}$
$=\mathbf{H}_{i}\mathbf{V}_{i}=\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]=\hat{\mathbf{H}}_{i}$
Here, $h$ represents the number of attention heads, and the input $\mathbf{x}\in\mathbb{R}^{(N,M)}$ is divided into $\mathbf{x}_{1},...,\mathbf{x}_{h}$ based on $h$. The parameters $\mathbf{W}_{Qi},\mathbf{W}_{Ki}$, and $\mathbf{W}_{Vi}$ correspond to $\mathbf{x}_{i}$. The whole process of MHA can be represented in Figure 3. Figure 3.a represnets that the input $\mathbf{x}$ is split into $\mathbf{x}_{1}\cdots\mathbf{x}_{8}$ based on the number of heads. Figure 3.b represents $\mathbf{x}_{i}\mathbf{W}_{Qi}=\mathbf{Q}_{i},\mathbf{x}_{i}\mathbf{W}_{Ki}=\mathbf{K}_{i},\mathbf{x}_{i}\mathbf{W}_{Vi}=\mathbf{V}_{i}$. Figure 3c, d, e represent the process of $\operatorname{softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{T}}{\sqrt{M}}\right)\mathbf{V}_{i}$ and we omit $\sqrt{M}$. Figure 3.f represents $\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}$. Figure 3.g represents the whole process by a matrix multiplication of $(\mathbf{W}^{\prime}_{HVO})^{T}\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, where $\mathbf{W}^{\prime}_{HVO}$ is generated based $\mathbf{H}_{1}\cdots\mathbf{H}_{8}$, $\mathbf{W}_{V1}\cdots\mathbf{W}_{V8}$ and $\mathbf{W}_{O}$. It means that we could represent the whole complex MHA into a matrix multiplication. Next, we give the proof of this process.
Our objective is to express the MHA as $(\mathbf{W}^{\prime})^{T}_{HVO}\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$. Prior to transforming the MHA into the matrix-vector form, we need to conduct a comprehensive analysis and clearly define the research object. In Eq. (7), $\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)$ describes an engineering process that requires mathematical representation. The learning process for the input $\mathbf{x}$ primarily consists of two parts: $\mathbf{V}_{1},\ldots,\mathbf{V}_{h}$ and $\mathbf{H}_{1},\ldots,\mathbf{H}_{h}$. $\mathbf{H}_{i}$ is derived based on the parameters $\mathbf{W}_{Qi}$ and $\mathbf{W}_{Ki}$, and can be directly regarded as parameters matrix now, a deep explanation is given in Section 3.5.
Figure 4.a represents the computation process of $\mathbf{H}_{1}\mathbf{V}_{1}=\hat{\mathbf{H}}_{1},\ldots,\mathbf{H}_{h}\mathbf{V}_{h}=\hat{\mathbf{H}}$, and in Figure 4.b, we convert it into the matrix-vector form $\mathbf{W}_{WH}^{\prime}\diamond\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{WH}^{\prime}$ is generated from $\mathbf{H}_{1},\dots,\mathbf{H}_{8}$ and $\mathbf{W}_{V1},\dots,\mathbf{W}_{V8}$. In Figure 4.c, we present a simple example, Figure 4.c.1 depicts $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$, while Figure 4.c.2 represents the matrix-vector format of Figure 4.c.1. More ditails could be found in Appendix B.2.
Figure 5 illustrates the parameter transformation scenario after incorporating $\mathbf{W}_{O}$. Figure 5.a demonstrates an example of $\mathring{\mathbf{H}}=\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}$, while Figure 5.b rewrites Figure 5.a as $\mathbf{H}^{\prime}\diamond\mathbf{W}_{O}^{\prime}=\mathring{\mathbf{H}}^{\prime}$. Consequently, based on Figures 4 and 5, the entire MHA can be expressed as $\mathbf{W}_{HVO}^{\prime}\diamond\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{HVO}^{\prime}=\mathbf{W}_{WH}^{T}\diamond\mathbf{W}_{O}^{\prime}$. So the whole MHA can be written as:
$\mathring{\mathbf{H}}$ $=(\mathbf{W}_{HVO}^{\prime})^{T}(\mathbf{x}_{i})^{\prime}$ (8)
Here, $\mathbf{x}_{i}\in\mathbb{R}^{(NM,1)}$ and $\mathring{\mathbf{H}}\in\mathbb{R}^{(NM,1)}$ are the input and output of $i$th layer, respectively, while $\mathbf{W}_{HVO}^{\prime}\in\mathbb{R}^{(NM,NM)}$ denotes matrices generated in accordance with $\mathbf{H}$, $\mathbf{W}_{V}$ and $\mathbf{W}_{O}$. In this manner, we have expressed MHA as a matrix-vector multiplication. This matrix multiplication representation provides a more concise way to express MHA mechanism.
Based on the derivations above and referring to Figure 4, we can gain a deeper understanding of the learning process in MHA. In particular, MHA can be regarded as a specialized form of linear transformation, which involves firstly dividing $\mathbf{x}$ into $\mathbf{x}_{1}...\mathbf{x}_{h}$ based on the number of heads, then transforming $\mathbf{x}_{i}$ into $\mathbf{x}_{i}^{\prime}$, and subsequently utilizing $\mathbf{H}_{i}^{\prime}$, $\mathbf{W}_{Vi}^{\prime}$, and $\mathbf{W}_{Oi}^{\prime}$ to generate the corresponding parameters $\mathbf{W}_{HVOi}^{\prime}$ for $\mathbf{x}_{i}^{\prime}$.
3.5 What makes Transformer to be the winner in LLMs?
The Transformer architecture is founded on two pivotal components: FFN and MHA. FFN is comprised of linear operations. In Sections 3.3 and 3.4, we have showcased the matrix-vector representations for both Linear and MHA. In this section, we delve into why the Transformer has emerged as a key model in the domain of LLMs.
From a mathematical perspective, both the MLP and the Transformer serve as concrete implementations of the UAT. Through the matrix-vector approach, it is straightforward to demonstrate that convolutional networks also realize the UAT, thus sharing the same foundational mathematical principles. However, the question arises: Why does the Transformer hold a dominant position in LLMs? The answer lies in the ingenious design of the MHA. Unlike traditional linear operations, the parameters of MHA are dynamic, adjusting in response to the input data. While the UAT endows models with a powerful approximation capability, fixing the parameters in linear operations means that the function that an MLP can approximate remains unchanged, as illustrated in Figure 6.a. In contrast, the Transformer can adjust its parameters by $\mathbf{H}_{i}^{\prime}$, based on the input. These parameters, $\mathbf{H}_{i}^{\prime}$, should be regarded as being generated in response to the input, as depicted in Figure 6.b. This adaptability implies that the approximated function can vary with the input. This characteristic also explains the strong generalization capability of Transformer-based neural networks, as they can approximate different functions tailored to the requirements of various tasks, such as those needed for language translation versus mathematical problem-solving.
Figure 6: The approximation ways of the MLP (a) and Transformer (b) networks.
4 Discussion
Leveranging on our proof in the previous section that the Transformer is the tangible embodiment of UAT, in this section we address the following critical problems and emplore the technical strategies of LLMs: Why does the Transformer architecture possess such power in enabling intelligent language models, such as translation and programming? What enables LLMs’ capacity for ICL? How does the LoRA scheme effectively fine-tune LLMs? What justifies the feasibility of pruning LLMs?
4.1 The fundamental reason for the power of Transformers in LLMs
Theoretically, we have established that Transformer networks act as concrete implementations of the UAT, leveraging its potent capability to approximate any Borel measurable function, thus ensuring the high-dimensional function approximation advantage of models based on UAT. While UAT possesses immense function approximation abilities, it inherently lacks the capacity of approximating multiple functions simultaneously. However, language tasks are by nature diverse, potentially requiring the approximation of different functions based on the input. For instance, summarizing, translating, and continuing the same text content, the function space remains entirely the same except for minor variations in prompt words. Without the ability to dynamically approximate functions based on input, but merely fitting based on a general function trend, it would lead to approximating the same function, hence producing identical or similar output. Therefore, LLMs need to distinguish and adapt to these nearly identical functions, thereby dynamically generating response functions based on input. The MHA mechanism in Transformers provides LLMs with the capability to dynamically approximate relevant functions based on input. In particular, MHA mechanism allows for the dynamic adjustment of UAT’s parameters according to the input, thereby laying the theoretical foundation for Transformer-based LLMs to simultaneously tackle multifaceted tasks such as translation, continuation, summarization, code generation, and solving mathematical problems.
4.2 What enables LLMs to possess ICL capability?
Contextual interaction, as the core capability of LLMs, permeates every phase from training and fine-tuning to prediction. ICL, multi-step reasoning, and instruction following are intuitive manifestations of this contextual interaction. Leveraging their context-sensitive interaction capabilities, LLMs can exhibit behaviors consistent with ICL, multi-step inference, and instruction following, which are tailored based on contextual cues. Furthermore, contemporary LLMs are evolving towards handling longer contexts; however, this advancement is accompanied by a dramatic increase in computational resource demands due to text expansion. A notable solution to this challenge is LongLoRA Chen et al. (2023), which fundamentally also utilizes contextual interaction mechanisms to facilitate training on extended sequences.
So, how does this contextual interaction capability arise within LLMs? The formula $\mathring{\mathbf{H}}=(\mathbf{W}_{HVO}^{\prime})^{T}(\mathbf{x}^{i})^{\prime}$ in Figure 5 reveals this mode of contextual interaction. Since $(\mathbf{W}_{HVO}^{\prime})^{T}$ represents a dense matrix (almost devoid of zero elements and whose internal elements are highly correlated), each element in $\mathring{\mathbf{H}}$ encapsulates comprehensive information from both preceding and subsequent contexts. This learning of holistic contextual information constitutes the foundation of contextual interaction within LLMs.
4.3 What justifies the feasibility of pruning LLMs?
Due to the massive size of parameters in LLMs and the subsequent high demand for computational resources, pruning LLMs is pivotal for their deployment. A legitimate question to ask is why LLMs are amenable to pruning? The rationale lies in the presence of excessively low-weight parameters in certain layers of LLMs. To understand this, we can directly analyze it from the perspective of the formula underlying the UAT:
$|\sum_{j=1}^{N}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)-f(\mathbf{x})|<\varepsilon\quad.$ (9)
for all $\mathbf{x}\in\mathbf{I}_{n}$. Let’s assume $\Lambda=\{1,2...N\},\Lambda_{1}\cup\Lambda_{2}=\Lambda,\Lambda_{1}\cap\Lambda_{2}=\emptyset$ and $|\sum_{j\in\Lambda_{1}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)|\to 0$. Then we have:
$|\sum_{j\in\Lambda_{1}}\alpha_{j}\sigma\left(\mathbf{W}j^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$ (10)
$+$ $\sum_{j\in\Lambda_{2}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)-f(\mathbf{x})|<\varepsilon$
Since $|\sum_{j\in\Lambda_{1}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)|\to 0$, we have the following inequality:
$\big{|}|\sum_{j\in\Lambda_{1}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)-f(\mathbf{x})\mid$ (11)
-- $|\sum_{j\in\Lambda_{2}}\alpha_{j}\sigma\left(\mathbf{W}j^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)|\big{|}<\varepsilon$
Therefore, we have:
$|\sum_{j=1}^{N_{1}}\alpha_{j}\sigma\left(\mathbf{W}j^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)-f(\mathbf{x})|$ (12)
$<\varepsilon+|\sum_{j=1}^{N_{2}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)|$
Hence, when parameters in certain layers are small enough, we can directly remove those layers since their impact on the final result is minimal.
Figure 7: The LoRA porcess (a), and the operation from the perspective of UAT (b).
Figure 8: The comparison between Human and LLMs.
4.4 How does the LoRA scheme effectively fine-tune LLMs?
Owing to the substantial computational resources required for training LLMs and their strong generalization capabilities, we believe the focus should be redirected towards efficiently leveraging pre-trained models. Repeatedly training models from scratch incurs considerable computational overhead, rendering the endeavor of repurposing well-trained models in novel tasks both practically valuable and resource-saving. A currently prominent solution addressing this issue is the LoRA methodology.
The LoRA scheme, illustrated in Figure 7.a, adopts a streamlined approach where $T_{i}$ symbolically represents the transformations within the network. Given our prior establishment that these transformations can be cast into matrix-vector operations, from the perspective of the UAT, LoRA fundamentally constitutes a layer-wise fine-tuning tailored to specific task characteristics. This fine-tuning mechanism is depicted in Figures 7.b.1 and Figure 7.b.2, where $\mathbf{W}^{\prime}_{amend}$ signifies the parameters post-refinement, thereby embodying an adaptive calibration of the UAT’s underlying parameters to enhance task-specific performance.
4.5 Rethinking LLMs
The capabilities of LLMs have become so formidable that their language processing prowess is increasingly approaching human levels, prompting a core question: How do LLMs differ from humans in language processing? Figure 8 elucidates the contrast between the patterns employed by LLMs and humans. Both start with language encoding—humans through a character-based system and LLMs via numerical arrays—essentially, there is little difference at this foundational level. Given the polysemy of words, context determination is crucial: humans understand context through the activation and transmission of neurons in the human brain, while LLMs use patterns of parameter activation to approximate corresponding functions. Here, the activation of network parameters parallels the neural electrical signalling in the brain, and the approximation of functions corresponds to human comprehension of semantics. From this perspective, the difference in language processing between humans and LLMs seems minimal.
We posit that the main differences lie in human autonomy and multimodal perception. Human autonomy enables self-directed learning, allowing knowledge to be validated within real-world contexts and personal experiences, whereas LLMs are confined to function fitting based on the majority of corpora in their datasets, with the goal of optimizing loss minimization. Human multimodality encompasses interpreting the world through senses such as vision, hearing, taste, and touch, infusing vocabulary with rich meanings, while LLMs are inherantly adept to to single-modality numerical representations.
Another significant issue in LLMs pertains to the segmentation of word embeddings within the MHA, as shown in Figure 4.a. Does this fragmentation still capture the full meaning of words comprehensively? If there are eight attention heads, does that mean a word has eight different encodings, suggesting a form of modality diversity, which then interacts through MHA to extract contextual insights? Furthermore, by considering LLMs as implementations of UAT, can we directly compute the weights necessary for function approximation from the data, thereby significantly reducing training overhead?
5 Conclusion
In this paper, we delve into the theoretical underpinnings of LLMs, demonstrating that contemporary LLMs, primarily constructed with Transformer architectures, embody concrete manifestations of the UAT. The remarkable generalization prowess exhibited by LLMs is attributed to their MHA modules, which enable the adaptation to and approximation of diverse functions based on the presented input data. Contextual interaction emerges as a paramount capability for LLMs, manifesting such abilities as ICL, instruction following, and contextual reasoning. These competencies are enabled by the Transformer’s innate capacity to learn from context.
Expanding upon this understanding, we have provided a rigorous theoretical grounding for key techniques employed in LLMs, including LoRA for efficient fine-tuning and pruning for model compression, elucidating their effectiveness through the lens of the UAT. By leveraging the theoretical framework provided by the UAT, not only can existing methodologies be explained but also avenues for the future evolution of LLMs are illuminated.
Limitation
Our work primarily delves into explaining a select few characteristics (ICL, Instruction following, and multi-step reasoning) and techniques (such as LoRA and Pruning) that we perceive as pivotal within current LLMs. While numerous advancements have been made to Transformer-based LLMs, their fundamental underpinnings largely align with the principles outlined herein. Nevertheless, due to space constraints, this paper does not aspire to exhaustively address every issue or attribute present across all LLMs. Instead, it focuses on elucidating those deemed most significant, thereby offering a concentrated insight into the core of LLMs. It is important to acknowledge that our scope, though carefully curated, is by no means exhaustive, and further exploration is warranted to gain a comprehensive understanding of the extensive landscape of LLMs enhancements and challenges.
Appendix A The Property of Diamond
[…]
$\mathbf{W}_{1}\diamond[\mathbf{x}\diamond\mathbf{W}_{2}]$ (13)
$=$ $\mathbf{W}_{1}\diamond[\mathbf{W}_{2}^{T}\mathbf{x}]$
$=$ $\mathbf{W}_{1}^{T}[\mathbf{W}_{2}^{T}\mathbf{x}]$
$=$ $\mathbf{W}_{1}^{T}\mathbf{W}_{2}^{T}\mathbf{x}=\mathbf{W}_{12}\mathbf{x}$
$=$ $(\mathbf{W}_{1}^{T}\mathbf{W}_{2}^{T})^{T}\diamond\mathbf{x}$
$=$ $(\mathbf{W}_{2}\mathbf{W}_{1})\diamond\mathbf{x}$
$=$ $\mathbf{W}_{2}\diamond\mathbf{W}_{1}\diamond\mathbf{x}$
$=$ $\mathbf{W}_{2}^{T}\diamond\mathbf{W}_{1}\diamond\mathbf{x}$
$[\mathbf{W}_{1}\diamond\mathbf{x}]\diamond\mathbf{W}_{2}$ (14)
$=$ $[\mathbf{W}_{1}^{T}\mathbf{x}]\diamond\mathbf{W}_{2}$
$=$ $\mathbf{W}_{2}^{T}[\mathbf{W}_{1}^{T}\mathbf{x}]$
$=$ $[\mathbf{W}_{2}^{T}\mathbf{W}_{1}^{T}]^{T}\diamond\mathbf{x}$
$=$ $[\mathbf{W}_{1}\mathbf{W}_{2}]\diamond\mathbf{x}$
$=$ $\mathbf{W}_{1}\diamond\mathbf{W}_{2}\diamond\mathbf{x}$
$=$ $\mathbf{W}_{1}^{T}\diamond\mathbf{W}_{2}\diamond\mathbf{x}$
Appendix B The Matrix-Vector Form of Linear and MHA
In this section, we will present the transformation process Linear and MHA into their corresponding matrix-vector forms.
B.1 The Matrix-Vector Form of Linear
Eq. ((15)) gives the transfomation method for Linear, it can be deduced that if the input data matrix has dimensions $(N,M)$ and the parameter matrix has dimensions $(N,N)$, the sparsity of the parameter matrix corresponding to the Linear operation is $1/M$.
$\left(\begin{array}[]{cccc}w_{1,1}&w_{1,2}&\cdots&w_{1,n}\\ w_{2,1}&w_{2,2}&\cdots&w_{2,n}\\ \vdots&\vdots&\vdots&\vdots\\ w_{n,1}&w_{n,2}&\cdots&w_{n,n}\end{array}\right)\left(\begin{array}[]{cccc}x_{1,1}&x_{1,2}&\cdots&x_{1,m}\\ x_{2,1}&x_{2,2}&\cdots&x_{2,m}\\ \vdots&\vdots&\vdots&\vdots\\ x_{n,1}&x_{n,2}&\cdots&x_{n,m}\end{array}\right)=\left(\begin{array}[]{cccc}y_{1,1}&y_{1,2}&\cdots&y_{1,m}\\ y_{2,1}&y_{2,2}&\cdots&y_{2,m}\\ \vdots&\vdots&\vdots&\vdots\\ y_{n,1}&y_{n,2}&\cdots&y_{n,m}\end{array}\right)$ (15)
$\rightarrow\left(\begin{array}[]{ccc}w_{1,1}&0&\cdots\\ 0&w_{1,1}&\cdots\\ \vdots&\vdots&\cdots\\ 0&0&\cdots\\ w_{1,2}&0&\cdots\\ 0&w_{1,2}&\cdots\\ \vdots&\vdots&\cdots\\ 0&0&\cdots\\ \vdots&\vdots&\cdots\\ w_{1,n}&0&\cdots\\ 0&w_{1,n}&\cdots\\ \vdots&\vdots&\cdots\\ 0&0&\cdots\end{array}\right)\diamond\left(\begin{array}[]{c}x_{1,1}\\ x_{1,2}\\ \vdots\\ x_{1,m}\\ x_{2,1}\\ x_{2,2}\\ \vdots\\ x_{2,m}\\ \vdots\\ x_{n,1}\\ x_{n,2}\\ \vdots\\ x_{n,m}\end{array}\right)=\left(\begin{array}[]{c}y_{1,1}\\ y_{1,2}\\ \vdots\\ y_{1,m}\\ y_{2,1}\\ y_{2,2}\\ \vdots\\ y_{2,m}\\ \vdots\\ y_{n,1}\\ y_{n,2}\\ \vdots\\ y_{n,m}\end{array}\right)$
B.2 The Matrix-Vector Form of MHA
Figure 1: The process of transforming $\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}}_{8})$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{WH}^{\prime}\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$.
Figure 1.a represents the computation process of obtaining $\mathbf{W}_{QKi}$ in the attention mechanism, while Figure 1.b computes $\mathbf{H}_{1},\dots,\mathbf{H}_{8}$ through the $softmax$ operation. Figure 1.c and Figure 1.e illustrates the process of converting $\operatorname{Concat}\left(\hat{\mathbf{H}}_{1}\mathbf{V}_{1},\ldots,\hat{\mathbf{H}}_{h}\mathbf{V}_{h}\right)$ into the matrix-vector form $\mathbf{W}_{WH}^{\prime}\diamond\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{WH}^{\prime}$ is generated from $\mathbf{H}_{1},\dots,\mathbf{H}_{8}$ and $\mathbf{W}_{V1},\dots,\mathbf{W}_{V8}$. Figure 1.c computes $\hat{\mathbf{H}}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$. In Figure 1.d, we provide a simple example demonstrating the conversion of $\hat{\mathbf{H}}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$ into $\mathbf{W}_{WHi}^{\prime}\mathbf{x}_{i}^{\prime}$. This process is divided into four parts: Figure 1.d.1 represents the general form of $\hat{\mathbf{H}}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$, while Figure 1.d.2 serves as a simple example of Figure 1.d.1. Figure 1.d.3 first rewrites Figure 1.d.2 using diamond multiplication as $\hat{\mathbf{H}}_{i}^{\prime}\diamond[\mathbf{x}_{i}^{\prime}\diamond\mathbf{W}_{Vi}^{\prime}]$, and then utilizes the property of diamond multiplication to express it as $[\mathbf{W}_{Vi}^{\prime T}\hat{\mathbf{H}}_{i}^{\prime}]\mathbf{x}_{i}^{\prime}$. Finally, Figure 1.d.4 is obtained as $\mathbf{W}_{WHi}^{\prime}\mathbf{x}_{i}^{\prime}$. It can be observed that $\mathbf{W}_{WHi}^{\prime}$ is a dense matrix. In Figure 1.e, we present the conversion of the entire $\mathbf{H}_{1}[\mathbf{x}_{1}\mathbf{W}_{V1}]...\mathbf{H}_{8}[\mathbf{x}_{8}\mathbf{W}_{V8}]$ into the matrix-vector form $\mathbf{W}_{WH}^{\prime}\mathbf{x}^{\prime}$. In Figure 1.f, we present a simple example of Figure 1.e. Figure 1.f.1 depicts $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$, while Figure 1.f.2 represents the Matrix-vector format of Figure 1.f.1.

### v3

@@
-Universal Approximation Theory: The basic theory for large language models
+Universal Approximation Theory: The Basic Theory for Transformer-based Large Language Models
 Abstract
-Language models have emerged as a critical area of focus in artificial intelligence, particularly with the introduction of groundbreaking innovations like ChatGPT.
-Large-scale Transformer networks have quickly become the leading approach for advancing natural language processing algorithms.
-Built on the Transformer architecture, these models enable interactions that closely mimic human communication and, equipped with extensive knowledge, can even assist in guiding human tasks.
-Despite their impressive capabilities and growing complexity, a key question remains – the theoretical foundations of large language models (LLMs).
-What is it about the Transformer architecture that makes it so effective for powering intelligent language applications, such as translation and coding?
-What underlies LLMs’ ability for In-Context Learning (ICL)?
-How does the LoRA scheme enhance the fine-tuning of LLMs?
-And what supports the practicality of pruning LLMs?
-To address these critical questions and explore the technological strategies within LLMs, we leverage the Universal Approximation Theory (UAT) to offer a theoretical backdrop, shedding light on the mechanisms that underpin these advancements.
-Universal Approximation Theory: The basic theory for large language models
-Wei Wang1, Qing Li1, 1The Hong Kong Polytechnic University, weiuat.wang@connect.polyu.hk, qing-prof.li@polyu.edu.hk
+Language models have become a focal point in the field of artificial intelligence, especially following the emergence of ChatGPT.
+Large-scale Transformer networks have quickly become the primary method for advancing natural language processing algorithms.
+These models, built on the Transformer architecture, are capable of engaging in near-human-like interactions and can even assist in guiding human tasks.
+Despite their impressive capabilities, there remain some critical theoretical questions in the development of large language models (LLMs): What makes the Transformer architecture so effective in driving intelligent language applications, such as translation and coding?
+What underlies the in-context learning (ICL) capability of LLMs?
+How does the LoRA approach enhance the fine-tuning of LLMs?
+What supports the feasibility of pruning in LLMs?
+To address these key questions and explore the technical strategies within LLMs, we draw on the Universal Approximation Theory (UAT) to provide a theoretical foundation, revealing the mechanisms behind these advancements.
 1 Introduction
@@
 Thus, endowing machines with the ability to understand and generate language marks a significant leap towards the realization of true artificial intelligence.
-The emergence of models such as the ChatGPT Radford and Narasimhan (2018); Brown et al.
-(2020); Achiam et al.
-(2023), the Llama Touvron et al.
-(2023), and the PaLM Chowdhery et al.
-(2023) vividly demonstrates this point.
-A distinctive feature of LLMs is their immense parameter size Achiam et al.
-(2023); Touvron et al.
-(2023); Chowdhery et al.
-(2023); Chen et al.
-(2021); Zeng et al.
-(2022), often amounting to hundreds of billions or even trillions (for instance, GPT-3’s Brown et al.
-(2020) 175 billion parameters and PaLM’s Chowdhery et al.
-(2023) 540 billion parameters).
-This vast parameter scale lays the foundation for their exceptional language processing capabilities and enables them to exhibit almost human-like traits, such as ICL  Brown et al.
-(2020); Dong et al.
-(2022), instruction following Sanh et al.
-(2021); Ouyang et al.
-(2022); Wei et al.
-(2021), and multi-step reasoning Wei et al.
-(2022).
-Notably, these colossal models are predominantly trained by tech giants like Google and Microsoft using large-scale GPU clusters Zhao et al.
-(2023), sparking a research fervor on how to efficiently fine-tune them with limited GPU resources.
-The advent of Lora Hu et al.
-(2021) fine-tuning technology has provided an effective pathway for this, allowing for the fine-tuning of large models under resource constraints without the need to adjust all parameters of the original model comprehensively.
-Moreover, model pruning techniques Sun et al.
-(2023); Ma et al.
-(2023) are crucial for deploying large models in resource-constrained environments, aiming to reduce the model size for operation on smaller devices.
-Faced with the challenge of processing long texts, such as generating summaries or answering questions based on extensive documents—which traditionally requires substantial computational resources—technologies like LongLora Chen et al.
-(2023) have been developed to tackle the difficulties of processing long contextual texts, further expanding the application boundaries of LLMs.
+The emergence of models such as the ChatGPT (Radford and Narasimhan 2018; Brown et al.
+2020; Achiam et al.
+2023), the Llama (Touvron et al.
+2023), and the PaLM (Chowdhery et al.
+2023) vividly demonstrates this point.
+A distinctive feature of LLMs is their immense parameter size (Achiam et al.
+2023; Touvron et al.
+2023; Chowdhery et al.
+2023; Chen et al.
+2021; Zeng et al.
+2022), often amounting to hundreds of billions or even trillions (for instance, GPT-3’s (Brown et al.
+2020) 175 billion parameters and PaLM’s (Chowdhery et al.
+2023) 540 billion parameters).
+This vast parameter scale lays the foundation for their exceptional language processing capabilities and enables them to exhibit almost human-like traits, such as ICL  (Brown et al.
+2020; Dong et al.
+2022), instruction following (Sanh et al.
+2021; Ouyang et al.
+2022; Wei et al.
+2021), and multi-step reasoning (Wei et al.
+2022).
+Notably, these colossal models are predominantly trained by tech giants like Google and Microsoft using large-scale GPU clusters (Zhao et al.
+2023), sparking a research fervor on how to efficiently fine-tune them with limited GPU resources.
+The advent of Lora (Hu et al.
+2021) fine-tuning technology has provided an effective pathway for this, allowing for the fine-tuning of large models under resource constraints without the need to adjust all parameters of the original model comprehensively.
+Moreover, model pruning techniques (Sun et al.
+2023; Ma, Fang, and Wang 2023) are crucial for deploying large models in resource-constrained environments, aiming to reduce the model size for operation on smaller devices.
 Today, LLMs possess a diverse range of functionalities, from translation and text summarization to automatic code generation, demonstrating their versatility.
-Despite LLMs being on a fast track of development towards higher intelligence and reliability, the theoretical foundation behind them remains largely unexplored, shrouded in clouds of uncertainty.
-The scientific community is actively seeking to uncover the inner mechanisms behind their powerful capabilities, including analyses of ICL mechanism Xie et al.
-(2021); Min et al.
-(2022), among others.
-Against this backdrop, we establish the UAT Cybenko (2007); Hornik et al.
-(1989) as the mathematical essence of LLMS, and utilize the UAT lens to clarify pivotal technologies and phenomena within contemporary LLMs, aiming to illuminate this central theoretical conundrum from a fresh perspective.
+Despite the rapid advancements of LLMs towards higher intelligence and reliability, the theoretical foundations behind these models remain largely unexplored and shrouded in uncertainty.
+The scientific community is actively seeking to uncover the intrinsic mechanisms behind their powerful capabilities, including the analysis of ICL mechanisms (Xie et al.
+2021; Min et al.
+2022).
+In this context, we establish the UAT (Cybenko 2007; Hornik, Stinchcombe, and White 1989) as the mathematical essence of Transformer-based LLMs and use the perspective of UAT to elucidate key technologies and phenomena within LLMs, aiming to provide explanations from a new angle.
+While previous studies have explored the relationship between Transformers and UAT (Yun et al.
+2019, 2020; Kratsios et al.
+2021; Alberti et al.
+2023), these studies are based on a series of complex assumptions and derivations, lacking generality and failing to intuitively address some of the issues present in LLMs.
+In contrast, this paper aims to provide proof that bridges Transformers and UAT, grounded in the foundational mathematical form of UAT as established by Cybenko; Hornik, Stinchcombe, and White.
+Our proof is characterized by its simplicity, absence of assumptions, and generality, making it applicable to Convolutional Neural Networks (CNNs) as well (Wang and Li 2024b).
+Our approach not only provides intuitive answers to some of the challenges within LLMs but also aids in the design of deep learning networks from a theoretical standpoint (Wang and Li 2024a).
+These capabilities are not addressed by previous theories.
 Our contributions are as follows:
@@
   • We explain the characteristics of LLMs, such as ICL, instruction following, multi-step reasoning, and the technologies applied within LLMs like Lora, pruning, and LLMs’ strong generalization capabilities.
-  • We offer insights into the future development of LLMs.
 Our article is structured as follows: In Section 2, we begin by introducing the UAT and propose that to demonstrate the Transformer’s adherence to UAT, it is necessary to show that both Linear and Multi-Head Attention (MHA) can be represented in the form of matrix-vector multiplication.
@@
 2 The Universal Approximation Theory
-Up to now, the UAT (Cybenko, 2007) is the most widely recognized fundamental theory in deep learning.
-However, this theorem is only applicable to the simplest form of neural networks, the multilayer perceptron (MLP) (Cybenko, 2007; Popescu et al., 2009).
+Up to now, the UAT (Cybenko 2007) is the most widely recognized fundamental theory in deep learning.
+However, this theorem is only applicable to the simplest form of neural networks, the multilayer perceptron (MLP) (Cybenko 2007; Popescu et al.
+2009).
 Due to the increased complexity of Transformer networks, they cannot be mathematically expressed in the same form as UAT.
@@
 This implies that, when $N$ is sufficiently large, a neural network can approximate any continuous function on a closed interval.
-Hornik et al.
-(1989) further demonstrates that multilayer feedforward networks also conform to the UAT, capable of approximating arbitrary Borel measurable functions.
+Hornik, Stinchcombe, and White (1989) further demonstrates that multilayer feedforward networks also conform to the UAT, capable of approximating arbitrary Borel measurable functions.
 Observing Equation (1), where the function $G(\mathbf{x})$ yields a scalar output in $\mathbb{R}$, the scenario expands naturally when $G(\mathbf{x})$ maps to $\mathbb{R}^{m}$, requiring the approximation in each dimension.
-It becomes evident that to accommodate this multidimensional output, a simple extension to Equation (1) suffices: the transformation matrix $\mathbf{W}_{j}$ is revised to the space $\mathbb{R}^{n\times m}$, the bias term $\theta_{j}$ is recast as a vector in $\mathbb{R}^{m}$, and $\alpha_{j}$ is reshaped into a diagonal vector.
-Nevertheless, in these demonstrations, the theorem does not straightforwardly apply to Transformer architectures.
-However, if we succeed in reformulating both the Linear ($\mathbf{x}|\mathbf{W}$) and MHA ($\mathbf{x}|\mathbf{W}$) components into a unified representation, $\mathbf{W}^{\prime}\mathbf{x}^{\prime}$, where $\mathbf{W}$ and $\mathbf{x}$ represent the parameters and inputs for each component respectively, and $\mathbf{W}^{\prime}$ is derived from $\mathbf{W}$ while $\mathbf{x}^{\prime}$ signifies a column vector (for details, refer to Section 3.1).
+It becomes evident that to accommodate this multidimensional output, a simple extension to Equation (1) suffices: the transformation matrix $\mathbf{W}_{j}$ is revised to the space $\mathbb{R}^{n\times m}$, the bias term $\theta_{j}$ is recast as a vector in $\mathbb{R}^{m}$, and $\alpha_{j}$ is reshaped into a matrix.
+Nevertheless, in these formulas, the theorem does not straightforwardly apply to Transformer architectures.
+However, if we succeed in reformulating both the Linear ($\mathbf{x}|\mathbf{W}$) and MHA ($\mathbf{x}|\mathbf{W}$) components into a unified representation, $\mathbf{W}^{\prime}\mathbf{x}^{\prime}$, where $\mathbf{W}$ and $\mathbf{x}$ represent the parameters and inputs for each component respectively, and $\mathbf{W}^{\prime}$ is derived from $\mathbf{W}$ while $\mathbf{x}^{\prime}$ signifies a column vector derived from $\mathbf{x}$ (for details, refer to Section 3.1).
 It is easy to prove that multi-layer Transformer networks are concrete embodiments of UAT.
@@
 In order to describe this transformation process clearly, we propose a new computational method called the Diamond Multiplication Method, denoted by the symbol $\diamond$.
+(The purpose of proposing this method is to clearly illustrate the calculation process in turning MHA into Matrix-Vector format.
+If we use direct matrix multiplication, some computational details cannot be visually distinguished.)
 $\left(\begin{array}[]{cccc}w_{1,1}&w_{1,2}&\cdots&w_{1,n}\\ w_{2,1}&w_{2,2}&&w_{2,n}\\ \vdots&\vdots&\vdots&\vdots\\ w_{m,1}&w_{m,2}&&w_{m,n}\end{array}\right)\diamond\left(\begin{array}[]{c}x_{1}\\ x_{2}\\ \vdots\\ x_{m}\end{array}\right)$ (4)
-$=\left(\begin{array}[]{c}w_{1}x_{1}+w_{31}*x_{2}+\cdots w_{m,1}x_{m}\\ w_{1,2}x_{1}+w_{2,2}x_{2}+\cdots w_{m,2}x_{m}\\ \vdots\\ w_{1,n}x_{1}+w_{2,n}x_{2}+\cdots w_{mn}x_{m}\end{array}\right)$
+$=\left(\begin{array}[]{c}w_{1}x_{1}+w_{21}x_{2}+\cdots w_{m,1}x_{m}\\ w_{1,2}x_{1}+w_{2,2}x_{2}+\cdots w_{m,2}x_{m}\\ \vdots\\ w_{1,n}x_{1}+w_{2,n}x_{2}+\cdots w_{mn}x_{m}\end{array}\right)$
 Let $\mathbf{W}\in\mathbb{R}^{(m,n)}$ and $\mathbf{x}\in\mathbb{R}^{(m,1)}$, the Diamond Matrix Multiplication is defined as $\mathbf{y}=\mathbf{W}\diamond\mathbf{x}$, where $\mathbf{y}\in\mathbb{R}^{(m,1)}$.
@@
 A detailed calculation process is provided in Eq.
-((4)).
+(4).
 The computation process for $\mathbf{y}=\mathbf{x}\diamond\mathbf{W}$ is identical to that of $\mathbf{W}\diamond\mathbf{x}$, implying that $\mathbf{W}\diamond\mathbf{x}=\mathbf{x}\diamond\mathbf{W}$.
-However, it should be noted that the Diamond Matrix Multiplication does not possess the properties of associativity and distributivity.
 Furthermore, based on its computation method, it can be inferred that the Diamond Matrix Multiplication is related to conventional matrix multiplication, specifically: $\mathbf{W}\diamond\mathbf{x}=\mathbf{W}^{T}\mathbf{x}$.
@@
 $\mathbf{x}_{i}^{\prime}$, $\mathbf{x}_{i+1}^{\prime}$, and $\mathbf{W}_{i}^{\prime}$ are generated based on $\mathbf{x}^{i}$, $\mathbf{x}^{i+1}$, and $\mathbf{W}_{i}$ using the Matrix-Vector Method.
-The derivation of the general form can be found in Appendix B.1.
+For convenience, we use $\mathbf{W}_{i}^{\prime}$ to represent $(\mathbf{W}_{i}^{\prime})^{T}$.
+Therefore, $\mathbf{x}_{i+1}^{\prime}=\mathbf{W}_{i}^{\prime}\mathbf{x}_{i}^{\prime}$.
+So the FFN in the Transformer could be written as:
+$FFN(\mathbf{x})$ $=\mathbf{W}_{2}\sigma(\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1})+\mathbf{b}_{2}$ (6)
+$\rightarrow\mathbf{W}_{2}^{\prime}\sigma(\mathbf{W}_{1}^{\prime}\mathbf{x}^{\prime}+\mathbf{b}_{1}^{\prime})+\mathbf{b}_{2}^{\prime}$
 Figure 3: The Transformer process.
-Figure 4: The process of transforming $\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}}_{8})$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{WH}^{\prime}\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$.
-Figure 5: This diagram illustrates the process of transforming $\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}})\mathbf{W}_{O}$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{WHO}^{\prime}\diamond\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, different colors in the diagram correspond to different variables.
+Figure 4: The process of transforming $\hat{\mathbf{H}}=\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}}_{8})$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{HV}^{\prime}\diamond\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$.
+Figure 5: This diagram illustrates the process of transforming $\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}})\mathbf{W}_{O}$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{HVO}^{\prime}\diamond\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, different colors in the diagram correspond to different variables.
 3.4 Matrix-Vector Method for MHA
@@
 $\mathring{\mathbf{H}}=\operatorname{MultiHead}(\mathbf{Q},\mathbf{K},\mathbf{V})$ (7)
-$=$ $\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}$
+$=$ $\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}=\hat{\mathbf{H}}\mathbf{W}_{O}$
 $\operatorname{Attention}(\mathbf{x}_{i}\mathbf{W}_{Qi}=\mathbf{Q}_{i},\mathbf{x}_{i}\mathbf{W}_{Ki}=\mathbf{K}_{i},$ (8)
@@
 The whole process of MHA can be represented in Figure 3.
-Figure 3.a represnets that the input $\mathbf{x}$ is split into $\mathbf{x}_{1}\cdots\mathbf{x}_{8}$ based on the number of heads.
+Figure 3.a represents that the input $\mathbf{x}$ is split into $\mathbf{x}_{1}\cdots\mathbf{x}_{8}$ based on the number of heads.
 Figure 3.b represents $\mathbf{x}_{i}\mathbf{W}_{Qi}=\mathbf{Q}_{i},\mathbf{x}_{i}\mathbf{W}_{Ki}=\mathbf{K}_{i},\mathbf{x}_{i}\mathbf{W}_{Vi}=\mathbf{V}_{i}$.
@@
 Figure 3.g represents the whole process by a matrix multiplication of $(\mathbf{W}^{\prime}_{HVO})^{T}\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, where $\mathbf{W}^{\prime}_{HVO}$ is generated based $\mathbf{H}_{1}\cdots\mathbf{H}_{8}$, $\mathbf{W}_{V1}\cdots\mathbf{W}_{V8}$ and $\mathbf{W}_{O}$.
-It means that we could represent the whole complex MHA into a matrix multiplication.
+This means that we could represent the whole complex MHA in matrix multiplication.
 Next, we give the proof of this process.
-Our objective is to express the MHA as $(\mathbf{W}^{\prime})^{T}_{HVO}\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$.
+Our objective is to express the MHA as $(\mathbf{W}^{\prime}_{HVO})^{T}\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$.
 Prior to transforming the MHA into the matrix-vector form, we need to conduct a comprehensive analysis and clearly define the research object.
 In Eq.
-(7), $\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)$ describes an engineering process that requires mathematical representation.
+(7), $\hat{\mathbf{H}}=\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)$ describes an engineering process that requires mathematical representation.
 The learning process for the input $\mathbf{x}$ primarily consists of two parts: $\mathbf{V}_{1},\ldots,\mathbf{V}_{h}$ and $\mathbf{H}_{1},\ldots,\mathbf{H}_{h}$.
-$\mathbf{H}_{i}$ is derived based on the parameters $\mathbf{W}_{Qi}$ and $\mathbf{W}_{Ki}$, and can be directly regarded as parameters matrix now, a deep explanation is given in Section 3.5.
-Figure 4.a represents the computation process of $\mathbf{H}_{1}\mathbf{V}_{1}=\hat{\mathbf{H}}_{1},\ldots,\mathbf{H}_{h}\mathbf{V}_{h}=\hat{\mathbf{H}}$, and in Figure 4.b, we convert it into the matrix-vector form $\mathbf{W}_{WH}^{\prime}\diamond\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{WH}^{\prime}$ is generated from $\mathbf{H}_{1},\dots,\mathbf{H}_{8}$ and $\mathbf{W}_{V1},\dots,\mathbf{W}_{V8}$.
-In Figure 4.c, we present a simple example, Figure 4.c.1 depicts $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$, while Figure 4.c.2 represents the matrix-vector format of Figure 4.c.1.
-More ditails could be found in Appendix B.2.
+$\mathbf{H}_{i}$ is derived based on the parameters $\mathbf{x}_{i}$, $\mathbf{W}_{Qi}$ and $\mathbf{W}_{Ki}$.
+Figure 4.a represents the computation process of $\mathbf{H}_{1}\mathbf{V}_{1}=\hat{\mathbf{H}}_{1},\ldots,\mathbf{H}_{h}\mathbf{V}_{h}=\hat{\mathbf{H}}_{h}$.
+In Figure 4.b, we present a simple example of Figure 4.a.
+Figure 4.b depicts $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}],i=1,2$.
+In Figure 4.c we convert Figure 4.b into the matrix-vector form $\mathbf{W}_{HV}^{\prime}\diamond\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{HV}^{\prime}$ is generated from $\mathbf{H}_{1},\mathbf{H}_{2}$ and $\mathbf{W}_{V1},\mathbf{W}_{V2}$.
+More details can be found in Appendix B.
 Figure 5 illustrates the parameter transformation scenario after incorporating $\mathbf{W}_{O}$.
-Figure 5.a demonstrates an example of $\mathring{\mathbf{H}}=\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}$, while Figure 5.b rewrites Figure 5.a as $\mathbf{H}^{\prime}\diamond\mathbf{W}_{O}^{\prime}=\mathring{\mathbf{H}}^{\prime}$.
-Consequently, based on Figures 4 and 5, the entire MHA can be expressed as $\mathbf{W}_{HVO}^{\prime}\diamond\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{HVO}^{\prime}=\mathbf{W}_{WH}^{T}\diamond\mathbf{W}_{O}^{\prime}$.
+Figure 5.a demonstrates an example of $\mathring{\mathbf{H}}=\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}$, while Figure 5.b rewrites Figure 5.a as $\hat{\mathbf{H}}^{\prime}\diamond\mathbf{W}_{O}^{\prime}=\mathring{\mathbf{H}}^{\prime}$.
+Consequently, based on Figures 4 and 5, the entire MHA can be expressed as $\mathbf{W}_{HVO}^{\prime}\diamond\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{HVO}^{\prime}=\mathbf{W}_{HV}^{T}\diamond\mathbf{W}_{O}^{\prime}$ (Based on Figures 5, we can get $\mathbf{W}_{HVO}^{\prime}$, but the resulting expression is too complex and not easy to understand, so we did not compute the final result.
+However, it is easy to analyze that as long as the variables in $\mathbf{W}_{HV}^{T}$ and $\mathbf{W}_{O}^{\prime}$ are non-zero, the result is essentially a dense matrix).
 So the whole MHA can be written as:
-$\mathring{\mathbf{H}}$ $=(\mathbf{W}_{HVO}^{\prime})^{T}(\mathbf{x}_{i})^{\prime}$ (8)
-Here, $\mathbf{x}_{i}\in\mathbb{R}^{(NM,1)}$ and $\mathring{\mathbf{H}}\in\mathbb{R}^{(NM,1)}$ are the input and output of $i$th layer, respectively, while $\mathbf{W}_{HVO}^{\prime}\in\mathbb{R}^{(NM,NM)}$ denotes matrices generated in accordance with $\mathbf{H}$, $\mathbf{W}_{V}$ and $\mathbf{W}_{O}$.
+$\mathring{\mathbf{H}}^{\prime}$ $=(\mathbf{W}_{HVO}^{\prime})^{T}\mathbf{x}^{\prime}$ (9)
+Here, $\mathbf{x}\in\mathbb{R}^{(NM,1)}$ and $\mathring{\mathbf{H}}\in\mathbb{R}^{(NM,1)}$ are the input and output, respectively, while $\mathbf{W}_{HVO}^{\prime}\in\mathbb{R}^{(NM,NM)}$ denotes matrices generated in accordance with $\mathbf{H}$, $\mathbf{W}_{V}$ and $\mathbf{W}_{O}$.
 In this manner, we have expressed MHA as a matrix-vector multiplication.
-This matrix multiplication representation provides a more concise way to express MHA mechanism.
-Based on the derivations above and referring to Figure 4, we can gain a deeper understanding of the learning process in MHA.
-In particular, MHA can be regarded as a specialized form of linear transformation, which involves firstly dividing $\mathbf{x}$ into $\mathbf{x}_{1}...\mathbf{x}_{h}$ based on the number of heads, then transforming $\mathbf{x}_{i}$ into $\mathbf{x}_{i}^{\prime}$, and subsequently utilizing $\mathbf{H}_{i}^{\prime}$, $\mathbf{W}_{Vi}^{\prime}$, and $\mathbf{W}_{Oi}^{\prime}$ to generate the corresponding parameters $\mathbf{W}_{HVOi}^{\prime}$ for $\mathbf{x}_{i}^{\prime}$.
-3.5 What makes Transformer to be the winner in LLMs?
+This matrix multiplication representation provides a more concise way to express the MHA mechanism.
+For convenience, we use $\mathbf{W}^{\prime}$ to represent $(\mathbf{W}_{HVO}^{\prime})^{T}$. The MHA operation could be simply written as:
+$\mathring{\mathbf{H}}^{\prime}=\mathbf{W}^{\prime}\mathbf{x}^{\prime}$ (10)
+3.5 The UAT Format of Multi-Layer Transformer
 The Transformer architecture is founded on two pivotal components: FFN and MHA.
-FFN is comprised of linear operations.
-In Sections 3.3 and 3.4, we have showcased the matrix-vector representations for both Linear and MHA.
-In this section, we delve into why the Transformer has emerged as a key model in the domain of LLMs.
-From a mathematical perspective, both the MLP and the Transformer serve as concrete implementations of the UAT.
-Through the matrix-vector approach, it is straightforward to demonstrate that convolutional networks also realize the UAT, thus sharing the same foundational mathematical principles.
-However, the question arises: Why does the Transformer hold a dominant position in LLMs?
-The answer lies in the ingenious design of the MHA.
-Unlike traditional linear operations, the parameters of MHA are dynamic, adjusting in response to the input data.
-While the UAT endows models with a powerful approximation capability, fixing the parameters in linear operations means that the function that an MLP can approximate remains unchanged, as illustrated in Figure 6.a.
-In contrast, the Transformer can adjust its parameters by $\mathbf{H}_{i}^{\prime}$, based on the input.
-These parameters, $\mathbf{H}_{i}^{\prime}$, should be regarded as being generated in response to the input, as depicted in Figure 6.b.
-This adaptability implies that the approximated function can vary with the input.
-This characteristic also explains the strong generalization capability of Transformer-based neural networks, as they can approximate different functions tailored to the requirements of various tasks, such as those needed for language translation versus mathematical problem-solving.
-Figure 6: The approximation ways of the MLP (a) and Transformer (b) networks.
+In Sections 3.3 and 3.4, we have showcased the matrix-vector representations for both FFN and MHA.
+In this section, we delve into why multi-layer Transformer is the implementation of UAT.
+Based on Equations 6 and 10, we can demonstrate that $i+1$-th layer ($i=0,1,2,\ldots$) Transformer is a concrete realization of the UAT, which can be written as (see Appendix C for more details):
+$\mathbf{x}_{i+1}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,1})+\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$ (11)
+where $\mathbf{x}_{i+1}$ is the output of ${i+1}$-th layer, $\mathbf{x}_{0}$ is the input of the network, $\mathbf{b}^{\prime}_{j,2}=(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})+\mathbf{W}^{\prime}_{j,2}UAT_{j-1}^{R}$, where $UAT_{j-1}^{R}=\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,3}\sigma(\mathbf{W}^{\prime}_{k,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,2})$.
+The term $\mathbf{b}^{\prime}_{j,2}$ is approximated by the $j$ layer of UAT with $\mathbf{x}_{0}$ as input.
+This enhances the model’s ability to dynamically adjust functions based on input.
+In the MHA mechanism, the parameters change dynamically with the input.
+Therefore, in the formula above, all $\mathbf{W}^{\prime}_{j,1}$, $\mathbf{W}^{\prime}_{j,2}$, and $\mathbf{W}^{\prime}_{j,3}$ parameters in layer $i$, where $j=1,\ldots,i$, are dynamically adjusted based on the input.
 4 Discussion
-Leveranging on our proof in the previous section that the Transformer is the tangible embodiment of UAT, in this section we address the following critical problems and emplore the technical strategies of LLMs: Why does the Transformer architecture possess such power in enabling intelligent language models, such as translation and programming?
+Leveraging on our proof in the previous section that the Transformer is the tangible embodiment of UAT, in this section we address the following critical problems and explore the technical strategies of LLMs: Why does the Transformer architecture possess such power in enabling intelligent language models, such as translation and programming?
 What enables LLMs’ capacity for ICL?
@@
 What justifies the feasibility of pruning LLMs?
-4.1 The fundamental reason for the power of Transformers in LLMs
-Theoretically, we have established that Transformer networks act as concrete implementations of the UAT, leveraging its potent capability to approximate any Borel measurable function, thus ensuring the high-dimensional function approximation advantage of models based on UAT.
-While UAT possesses immense function approximation abilities, it inherently lacks the capacity of approximating multiple functions simultaneously.
-However, language tasks are by nature diverse, potentially requiring the approximation of different functions based on the input.
-For instance, summarizing, translating, and continuing the same text content, the function space remains entirely the same except for minor variations in prompt words.
-Without the ability to dynamically approximate functions based on input, but merely fitting based on a general function trend, it would lead to approximating the same function, hence producing identical or similar output.
-Therefore, LLMs need to distinguish and adapt to these nearly identical functions, thereby dynamically generating response functions based on input.
-The MHA mechanism in Transformers provides LLMs with the capability to dynamically approximate relevant functions based on input.
-In particular, MHA mechanism allows for the dynamic adjustment of UAT’s parameters according to the input, thereby laying the theoretical foundation for Transformer-based LLMs to simultaneously tackle multifaceted tasks such as translation, continuation, summarization, code generation, and solving mathematical problems.
+4.1 What Makes Transformer So Powerful in LLMs?
+Theoretically, we have established that Transformer networks are concrete implementations of the UAT, enabling them to approximate any Borel measurable function.
+While UAT provides powerful function approximation capabilities, it inherently lacks the ability to approximate multiple functions simultaneously.
+However, language tasks are inherently diverse, often requiring the approximation of different functions based on the input.
+For instance, when summarizing, translating, or continuing the same text, the input functions are nearly identical, with only minor variations in the prompt.
+Without the ability to dynamically approximate functions based on the input, simply fitting a general function trend based on input will result in identical or similar outputs.
+To address this, LLMs must distinguish and adapt to these nearly identical functions, dynamically generating response functions based on the input.
+The MHA and residual mechanisms in Transformers equip LLMs with the ability to dynamically approximate relevant functions according to the input.
+Specifically, MHA allows for the dynamic adjustment of the weight parameters in UAT in response to the input, while the residual mechanism supports the dynamic approximation of bias terms, as illustrated in Figure 6.
+This theoretical foundation enables Transformer-based LLMs to handle a wide range of tasks, including translation, continuation, summarization, code generation, and solving mathematical problems.
+Figure 6: This diagram vividly demonstrates the output differences between an MLP and a Transformer when given the same input content (represented by the blue line in a) with only slight variations (represented by the black line in a).
+Since the input content is nearly identical and the MLP’s parameters are fixed, the MLP tends to produce the same output (b), influenced by the overall input function trend.
+However, the Transformer, with its ability to dynamically adapt to the input, generates different outputs based on the input variations.
+In the context of LLMs, the blue line in a can be understood as a consistent piece of text, while the black line represents different prompts, such as translation, continuation, or summarization.
+Furthermore, the MHA mechanism can capture global information (as shown in Figure 5, where each element in the output $\mathring{\mathbf{H}}^{\prime}$ contains global information), which helps generate content that is consistent with the context.
+This is crucial for language understanding that requires extensive contextual information.
+In contrast, 1D convolution employs a sliding convolution learning strategy, where the learned content is somewhat influenced by the size of the convolution kernel, leading to a more fragmented learning approach (for more information, please refer to Appendix D).
 4.2 What Enables LLMs to Possess ICL Capability?
@@
 Leveraging their context-sensitive interaction capabilities, LLMs can exhibit behaviors consistent with ICL, multi-step inference, and instruction following, which are tailored based on contextual cues.
-Furthermore, contemporary LLMs are evolving towards handling longer contexts; however, this advancement is accompanied by a dramatic increase in computational resource demands due to text expansion.
-A notable solution to this challenge is LongLoRA Chen et al.
-(2023), which fundamentally also utilizes contextual interaction mechanisms to facilitate training on extended sequences.
 So, how does this contextual interaction capability arise within LLMs?
-The formula $\mathring{\mathbf{H}}=(\mathbf{W}_{HVO}^{\prime})^{T}(\mathbf{x}^{i})^{\prime}$ in Figure 5 reveals this mode of contextual interaction.
-Since $(\mathbf{W}_{HVO}^{\prime})^{T}$ represents a dense matrix (almost devoid of zero elements and whose internal elements are highly correlated), each element in $\mathring{\mathbf{H}}$ encapsulates comprehensive information from both preceding and subsequent contexts.
+The formula $\mathring{\mathbf{H}}=\mathbf{W}^{\prime}\mathbf{x}^{\prime}$ in Figure 5 reveals this mode of contextual interaction.
+Since $\mathbf{W}^{\prime}$ represents a dense matrix (almost devoid of zero elements and whose internal elements are highly correlated), each element in $\mathring{\mathbf{H}}$ encapsulates comprehensive information from both preceding and subsequent contexts.
 This learning of holistic contextual information constitutes the foundation of contextual interaction within LLMs.
+(See Appendix D for more details)
 4.3 What Justifies the Feasibility of Pruning LLMs?
 Due to the massive size of parameters in LLMs and the subsequent high demand for computational resources, pruning LLMs is pivotal for their deployment.
-A legitimate question to ask is why LLMs are amenable to pruning?
+A legitimate question to ask is why LLMs are amenable to pruning.
 The rationale lies in the presence of excessively low-weight parameters in certain layers of LLMs.
@@
 Hence, when parameters in certain layers are small enough, we can directly remove those layers since their impact on the final result is minimal.
-Figure 7: The LoRA porcess (a), and the operation from the perspective of UAT (b).
+4.4 How Does the LoRA Scheme Effectively Fine-tune LLMs?
+Given the substantial computational resources required to train LLMs and their powerful generalization abilities, we believe that more efficient use of pre-trained models is essential.
+Re-training models from scratch incurs significant computational costs, so reusing well-trained models for new tasks is both practical and resource-efficient.
+A prominent solution to this challenge is the LoRA (Hu et al.
+2021), which can be expressed as follows:
+$\mathbf{h}=\mathbf{W}_{0}\mathbf{x}+\Delta\mathbf{W}\mathbf{x}$ (16)
+According to Eq. 11, we use LoRA to fine-tune the Linear operation in FFN can be written as:
+$\mathbf{x}_{i+1}$ $=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,1})+\sum_{j=1}^{i+1}(\mathbf{W}^{\prime}_{j,3}+\Delta\mathbf{W}_{j+1,3})$ (17)
+$\sigma((\mathbf{W}^{\prime}_{j,2}+\Delta\mathbf{W}_{j+1,2})\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$
+From Eq. 17, it can be seen that LoRA essentially fine-tunes the UAT parameters layer by layer for a specific task.
 Figure 7: The comparison between Human and LLMs.
-4.4 How does the LoRA scheme effectively fine-tune LLMs?
-Owing to the substantial computational resources required for training LLMs and their strong generalization capabilities, we believe the focus should be redirected towards efficiently leveraging pre-trained models.
-Repeatedly training models from scratch incurs considerable computational overhead, rendering the endeavor of repurposing well-trained models in novel tasks both practically valuable and resource-saving.
-A currently prominent solution addressing this issue is the LoRA methodology.
-The LoRA scheme, illustrated in Figure 7.a, adopts a streamlined approach where $T_{i}$ symbolically represents the transformations within the network.
-Given our prior establishment that these transformations can be cast into matrix-vector operations, from the perspective of the UAT, LoRA fundamentally constitutes a layer-wise fine-tuning tailored to specific task characteristics.
-This fine-tuning mechanism is depicted in Figures 7.b.1 and Figure 7.b.2, where $\mathbf{W}^{\prime}_{amend}$ signifies the parameters post-refinement, thereby embodying an adaptive calibration of the UAT’s underlying parameters to enhance task-specific performance.
 4.5 Rethinking LLMs
-The capabilities of LLMs have become so formidable that their language processing prowess is increasingly approaching human levels, prompting a core question: How do LLMs differ from humans in language processing?
-Figure 8 elucidates the contrast between the patterns employed by LLMs and humans.
-Both start with language encoding—humans through a character-based system and LLMs via numerical arrays—essentially, there is little difference at this foundational level.
-Given the polysemy of words, context determination is crucial: humans understand context through the activation and transmission of neurons in the human brain, while LLMs use patterns of parameter activation to approximate corresponding functions.
-Here, the activation of network parameters parallels the neural electrical signalling in the brain, and the approximation of functions corresponds to human comprehension of semantics.
-From this perspective, the difference in language processing between humans and LLMs seems minimal.
-We posit that the main differences lie in human autonomy and multimodal perception.
-Human autonomy enables self-directed learning, allowing knowledge to be validated within real-world contexts and personal experiences, whereas LLMs are confined to function fitting based on the majority of corpora in their datasets, with the goal of optimizing loss minimization.
-Human multimodality encompasses interpreting the world through senses such as vision, hearing, taste, and touch, infusing vocabulary with rich meanings, while LLMs are inherantly adept to to single-modality numerical representations.
-Another significant issue in LLMs pertains to the segmentation of word embeddings within the MHA, as shown in Figure 4.a.
-Does this fragmentation still capture the full meaning of words comprehensively?
-If there are eight attention heads, does that mean a word has eight different encodings, suggesting a form of modality diversity, which then interacts through MHA to extract contextual insights?
-Furthermore, by considering LLMs as implementations of UAT, can we directly compute the weights necessary for function approximation from the data, thereby significantly reducing training overhead?
+The capabilities of LLMs have become so advanced that their language processing abilities are approaching human levels, raising a core question: How do LLMs differ from humans in language processing?
+Figure 7 illustrates the comparison between the language processing processes of LLMs and humans.
+Both start with language encoding—humans encode language through a character-based system, while LLMs use numerical arrays.
+At this level, there is almost no difference.
+Given the ambiguity of words, determining context is crucial: humans understand context through the activation and transmission of neurons in the brain, while LLMs approximate the corresponding functions using UAT.
+Here, the input and output of network layers are analogous to the transmission of neural signals in the brain, and the function approximation corresponds to the final output of humans.
+From this perspective, the differences between humans and LLMs in language processing seem minimal.
+So how do we explain human understanding, analysis of language, and memory retrieval?
+Are these also capabilities of LLMs?
+First, we can consider the human brain as a combination of one or more UAT models, which are randomly initialized at birth.
+What supports such an assumption?
+The original neural network, the perceptron (Rosenblatt 1963), was designed based on the human neuron.
+Over time, as deep learning networks evolved, the difference in form between deep networks and the perceptron grew, leading people to no longer associate neural networks with human neurons.
+However, through our derivation, the Transformer can also be understood as an implementation of UAT, and the perceptron can be seen as a layer of UAT.
+The human brain can be considered as a cluster of multi-layer perception.
+Learning language in a social environment is akin to training the UATs in the brain.
+Therefore, what we call understanding, analysis, and memory retrieval are essentially processes of fitting outputs based on inputs.
+Understanding and analysis: When we respond to a question posed by someone, it can be seen as a simple example of understanding and analysis, which both humans and LLMs can accomplish.
+Humans may guess the answer to unfamiliar questions based on their experiences, which can be either correct or incorrect, and LLMs exhibit similar behavior.
+Memory retrieval: This is essentially fitting specific results based on particular words, as there is no actual database in the brain.
+For instance, recalling the experience of eating an apple for the first time in childhood is the brain fitting specific results based on those words; without those words, the memory would not surface.
+This is because the brain dynamically outputs results based on inputs and fills in the details of the event based on learned natural rules (We provide some examples about understanding, analysis of language, and memory retrieval in Appendix E).
+However, as humans grow, the weights in the brain are constantly updated, leading to potential memory distortions.
+Therefore, it is entirely reasonable for LLMs to make errors or generate hallucinations—these are just outputs produced based on existing weights and inputs, a problem humans also face.
+We believe the greatest advantage humans have over LLMs is their powerful multimodal and multitask processing abilities (which can be understood as the coordination of multiple UAT models to produce reasonable results).
+Another advantage is the interaction with the real world, which allows us to verify the knowledge we’ve learned in reality, enabling the brain to continuously optimize its parameters based on inputs.
+LLMs, on the other hand, are limited to function approximation within the corpora data.
 5 Conclusion
 In this paper, we delve into the theoretical underpinnings of LLMs, demonstrating that contemporary LLMs, primarily constructed with Transformer architectures, embody concrete manifestations of the UAT.
-The remarkable generalization prowess exhibited by LLMs is attributed to their MHA modules, which enable the adaptation to and approximation of diverse functions based on the presented input data.
+The remarkable generalization prowess exhibited by LLMs is attributed to their MHA modules and residual operation, which enable the adaptation to approximate diverse functions based on the presented input data.
 Contextual interaction emerges as a paramount capability for LLMs, manifesting such abilities as ICL, instruction following, and contextual reasoning.
-These competencies are enabled by the Transformer’s innate capacity to learn from context.
+These competencies are enabled by the Transformer’s MHA to learn from context.
 Expanding upon this understanding, we have provided a rigorous theoretical grounding for key techniques employed in LLMs, including LoRA for efficient fine-tuning and pruning for model compression, elucidating their effectiveness through the lens of the UAT.
 By leveraging the theoretical framework provided by the UAT, not only can existing methodologies be explained but also avenues for the future evolution of LLMs are illuminated.
-Limitation
-Our work primarily delves into explaining a select few characteristics (ICL, Instruction following, and multi-step reasoning) and techniques (such as LoRA and Pruning) that we perceive as pivotal within current LLMs.
-While numerous advancements have been made to Transformer-based LLMs, their fundamental underpinnings largely align with the principles outlined herein.
-Nevertheless, due to space constraints, this paper does not aspire to exhaustively address every issue or attribute present across all LLMs.
-Instead, it focuses on elucidating those deemed most significant, thereby offering a concentrated insight into the core of LLMs.
-It is important to acknowledge that our scope, though carefully curated, is by no means exhaustive, and further exploration is warranted to gain a comprehensive understanding of the extensive landscape of LLMs enhancements and challenges.
 Appendix A The Property of Diamond
@@
 $=$ $(\mathbf{W}_{2}\mathbf{W}_{1})\diamond\mathbf{x}$
-$=$ $\mathbf{W}_{2}\diamond\mathbf{W}_{1}\diamond\mathbf{x}$
 $=$ $\mathbf{W}_{2}^{T}\diamond\mathbf{W}_{1}\diamond\mathbf{x}$
@@
 $=$ $[\mathbf{W}_{1}\mathbf{W}_{2}]\diamond\mathbf{x}$
-$=$ $\mathbf{W}_{1}\diamond\mathbf{W}_{2}\diamond\mathbf{x}$
 $=$ $\mathbf{W}_{1}^{T}\diamond\mathbf{W}_{2}\diamond\mathbf{x}$
-Appendix B The Matrix-Vector Form of Linear and MHA
-In this section, we will present the transformation process Linear and MHA into their corresponding matrix-vector forms.
-B.1 The Matrix-Vector Form of Linear
+Figure 1: The process of transforming $\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}}_{8})$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{HV}^{\prime}\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$.
+Figure 2: The process of Transformer.
+Appendix B The Matrix-Vector Form of MHA
+In this section, we will present the transformation process of MHA into their corresponding matrix-vector forms.
+Figure 1.a represents the computation process of obtaining $\mathbf{W}_{QKi}$ in the attention mechanism, while Figure 1.b computes $\mathbf{H}_{1},\dots,\mathbf{H}_{8}$ through the softmax operation.
+Figure 1.c and Figure 1.e illustrates the process of converting $\operatorname{Concat}\left(\mathbf{H}_{1}\mathbf{V}_{1},\ldots,\mathbf{H}_{h}\mathbf{V}_{h}\right)$ into the matrix-vector form $\mathbf{W}_{HV}^{\prime}\diamond\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{HV}^{\prime}$ is generated from $\mathbf{H}_{1},\dots,\mathbf{H}_{8}$ and $\mathbf{W}_{V1},\dots,\mathbf{W}_{V8}$.
+Figure 1.c shows $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$.
+In Figure 1.d, we provide a simple example demonstrating the conversion of $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$ into $\mathbf{W}_{HVi}^{\prime}\mathbf{x}_{i}^{\prime}$.
+This process is divided into four parts: Figure 1.d.1 represents the general form of $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$, while Figure 1.d.2 serves as a simple example of Figure 1.d.1.
+Figure 1.d.3 first rewrites Figure 1.d.2 using diamond multiplication as $\mathbf{H}_{i}^{\prime}\diamond[\mathbf{x}_{i}^{\prime}\diamond\mathbf{W}_{Vi}^{\prime}]$, and then utilizes the property of diamond multiplication to express it as $(\mathbf{W}_{Vi}^{\prime})^{T}\diamond\mathbf{H}_{i}^{\prime}\diamond\mathbf{x}_{i}^{\prime}$ in Figure 1.d.4.
+Finally, Figure 1.d.5 is obtained as $\mathbf{W}_{HVi}^{\prime}\diamond\mathbf{x}_{i}^{\prime}=\mathbf{H}_{i}^{\prime}$.
+It can be observed that $\mathbf{W}_{HVi}^{\prime}$ is a dense matrix.
+In Figure 1.e, we present a simple example of $\mathbf{H}_{1}[\mathbf{x}_{1}\mathbf{W}_{V1}],\mathbf{H}_{2}[\mathbf{x}_{2}\mathbf{W}_{V2}]...$ into the matrix-vector form $\mathbf{W}_{HV}^{\prime}\mathbf{x}^{\prime}$.
+Figure 1.e.1 depicts $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$, while Figure 1.e.2 represents the Matrix-vector format of Figure 1.e.1.
+Appendix C The UAT Format of Transformer
+In this section, we will prove multi-layer Transformer is the implementation of UAT.
+C.1 The Properties of UAT
+Before expressing Transformers in the UAT format, we present two lemma regarding UAT.
+There are two cases for UAT-approximated functions: $f(\mathbf{x})\in\mathbb{R}$ and $f(\mathbf{x})\in\mathbb{R}^{m}$.
+The proof for the case where $f(\mathbf{x})\in\mathbb{R}$ can be inferred from $f(\mathbf{x})\in\mathbb{R}^{m}$.
+Therefore, we will only provide the proof for approximating $f(\mathbf{x})\in\mathbb{R}^{m}$ using UAT.
+Lemma 1. The mathematical form of UAT remains unchanged when multiplied by a matrix (constant).
+$G(\mathbf{x})$ $=\mathbf{\beta}\sum_{j=1}^{N}\mathbf{\alpha}_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\mathbf{\theta}_{j}\right)$ (3)
+$=\sum_{j=1}^{N}\mathbf{\beta}\mathbf{\alpha}_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\mathbf{\theta}_{j}\right)$
 Eq.
-((15)) gives the transfomation method for Linear, it can be deduced that if the input data matrix has dimensions $(N,M)$ and the parameter matrix has dimensions $(N,N)$, the sparsity of the parameter matrix corresponding to the Linear operation is $1/M$.
-$\left(\begin{array}[]{cccc}w_{1,1}&w_{1,2}&\cdots&w_{1,n}\\ w_{2,1}&w_{2,2}&\cdots&w_{2,n}\\ \vdots&\vdots&\vdots&\vdots\\ w_{n,1}&w_{n,2}&\cdots&w_{n,n}\end{array}\right)\left(\begin{array}[]{cccc}x_{1,1}&x_{1,2}&\cdots&x_{1,m}\\ x_{2,1}&x_{2,2}&\cdots&x_{2,m}\\ \vdots&\vdots&\vdots&\vdots\\ x_{n,1}&x_{n,2}&\cdots&x_{n,m}\end{array}\right)=\left(\begin{array}[]{cccc}y_{1,1}&y_{1,2}&\cdots&y_{1,m}\\ y_{2,1}&y_{2,2}&\cdots&y_{2,m}\\ \vdots&\vdots&\vdots&\vdots\\ y_{n,1}&y_{n,2}&\cdots&y_{n,m}\end{array}\right)$ (15)
-$\rightarrow\left(\begin{array}[]{ccc}w_{1,1}&0&\cdots\\ 0&w_{1,1}&\cdots\\ \vdots&\vdots&\cdots\\ 0&0&\cdots\\ w_{1,2}&0&\cdots\\ 0&w_{1,2}&\cdots\\ \vdots&\vdots&\cdots\\ 0&0&\cdots\\ \vdots&\vdots&\cdots\\ w_{1,n}&0&\cdots\\ 0&w_{1,n}&\cdots\\ \vdots&\vdots&\cdots\\ 0&0&\cdots\end{array}\right)\diamond\left(\begin{array}[]{c}x_{1,1}\\ x_{1,2}\\ \vdots\\ x_{1,m}\\ x_{2,1}\\ x_{2,2}\\ \vdots\\ x_{2,m}\\ \vdots\\ x_{n,1}\\ x_{n,2}\\ \vdots\\ x_{n,m}\end{array}\right)=\left(\begin{array}[]{c}y_{1,1}\\ y_{1,2}\\ \vdots\\ y_{1,m}\\ y_{2,1}\\ y_{2,2}\\ \vdots\\ y_{2,m}\\ \vdots\\ y_{n,1}\\ y_{n,2}\\ \vdots\\ y_{n,m}\end{array}\right)$
-B.2 The Matrix-Vector Form of MHA
-Figure 1: The process of transforming $\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}}_{8})$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{WH}^{\prime}\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$.
-Figure 1.a represents the computation process of obtaining $\mathbf{W}_{QKi}$ in the attention mechanism, while Figure 1.b computes $\mathbf{H}_{1},\dots,\mathbf{H}_{8}$ through the $softmax$ operation.
-Figure 1.c and Figure 1.e illustrates the process of converting $\operatorname{Concat}\left(\hat{\mathbf{H}}_{1}\mathbf{V}_{1},\ldots,\hat{\mathbf{H}}_{h}\mathbf{V}_{h}\right)$ into the matrix-vector form $\mathbf{W}_{WH}^{\prime}\diamond\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{WH}^{\prime}$ is generated from $\mathbf{H}_{1},\dots,\mathbf{H}_{8}$ and $\mathbf{W}_{V1},\dots,\mathbf{W}_{V8}$.
-Figure 1.c computes $\hat{\mathbf{H}}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$.
-In Figure 1.d, we provide a simple example demonstrating the conversion of $\hat{\mathbf{H}}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$ into $\mathbf{W}_{WHi}^{\prime}\mathbf{x}_{i}^{\prime}$.
-This process is divided into four parts: Figure 1.d.1 represents the general form of $\hat{\mathbf{H}}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$, while Figure 1.d.2 serves as a simple example of Figure 1.d.1.
-Figure 1.d.3 first rewrites Figure 1.d.2 using diamond multiplication as $\hat{\mathbf{H}}_{i}^{\prime}\diamond[\mathbf{x}_{i}^{\prime}\diamond\mathbf{W}_{Vi}^{\prime}]$, and then utilizes the property of diamond multiplication to express it as $[\mathbf{W}_{Vi}^{\prime T}\hat{\mathbf{H}}_{i}^{\prime}]\mathbf{x}_{i}^{\prime}$.
-Finally, Figure 1.d.4 is obtained as $\mathbf{W}_{WHi}^{\prime}\mathbf{x}_{i}^{\prime}$.
-It can be observed that $\mathbf{W}_{WHi}^{\prime}$ is a dense matrix.
-In Figure 1.e, we present the conversion of the entire $\mathbf{H}_{1}[\mathbf{x}_{1}\mathbf{W}_{V1}]...\mathbf{H}_{8}[\mathbf{x}_{8}\mathbf{W}_{V8}]$ into the matrix-vector form $\mathbf{W}_{WH}^{\prime}\mathbf{x}^{\prime}$.
-In Figure 1.f, we present a simple example of Figure 1.e.
-Figure 1.f.1 depicts $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$, while Figure 1.f.2 represents the Matrix-vector format of Figure 1.f.1.
+(3) shows the representation of UAT multiplying a matrix.
+Let $\mathbf{\alpha}_{j}=\mathbf{\beta}\mathbf{\alpha}_{j}$, and the result remains consistent with the original UAT mathematical form.
+Thus, it is proven that the mathematical form of UAT remains unchanged when multiplied by a matrix (constant).
+Lemma 2.
+Let the general term of the network be written as: $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, $i=0,1,2,3,\ldots$.
+The multi-layer network of this form still corresponds to the mathematical form of UAT, and we refer to such a network as a residual network.
+To prove that a multi-layer network with the general term written as $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$ corresponds to the UAT mathematical form, we first provide the forms of single-layer and two-layer networks, as shown in Eq.
+(4) and (5).
+$\mathbf{x}^{\prime}_{1}=(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ (4)
+$\mathbf{x}^{\prime}_{2}$ $=\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{1}+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{2,2})+\mathbf{b}^{\prime}_{2,3}$ (5)
+$=(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{2,3})+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{2,2})$
+=$=\{\mathbf{W}^{\prime}_{2,1}[(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})]+\mathbf{b}^{\prime}_{2,3}\}$
+$+\mathbf{W}^{\prime}_{2,3}\sigma\{\mathbf{W}^{\prime}_{2,2}[(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})]+\mathbf{b}^{\prime}_{2,2}\}$
+=$=\{\mathbf{W}^{\prime}_{2,1}(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{b}^{\prime}_{2,3}+\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})\}$
+$+\mathbf{W}^{\prime}_{2,3}\sigma\{\mathbf{W}^{\prime}_{2,2}({\mathbf{W}}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{b}^{\prime}_{2,2}+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})\}$
+=$=\{(\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,1}}\mathbf{x}^{\prime}_{0}+\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,3}})+\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})\}$
+$+\mathbf{W}^{\prime}_{2,3}\sigma\{\underline{\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}}\mathbf{x}^{\prime}_{0}+\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}\}$
+In Eq.
+(5), let $\mathbf{W}^{\prime}_{2,1}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,1}}$, $\mathbf{b}^{\prime}_{2,3}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,3}}$, $\mathbf{W}^{\prime}_{1,3}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}}$, $\mathbf{W}^{\prime}_{2,3}=\underline{\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}}$, and $\mathbf{b}^{\prime}_{2,2}=\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}$.
+Thus, Eq.
+(6) can be written as:
+$\mathbf{x}^{\prime}_{2}$ $=(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,1})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ (6)
+$+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,2})$
+Why can we combine the above terms?
+Once the network training is complete, we know these parameters, so they can be directly calculated and thus combined.
+And $\mathbf{b}^{\prime}_{2,2}=\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}$ clearly has the same mathematical form as UAT.
+We can understand it as dynamically fitting the bias term $\mathbf{b}^{\prime}_{2,2}$ using UAT based on the input.
+Therefore, we have proven that the mathematical form of one and two-layer residual networks is consistent with the UAT mathematical form.
+Assume that the mathematical form of the first $i$ layers of the residual network is consistent with UAT.
+Our goal is to prove that the mathematical form of the $i+1$-th layer of the residual network is still consistent with UAT.
+For the convenience of expression, we make the following definition: since the mathematical form of the first $i$ layers of the residual network is consistent with UAT, we write $\mathbf{x}^{\prime}_{i}$ as $\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,3})+UAT^{R}_{i}$, where the first term is written separately and the rest is written as the remainder term $UAT^{R}_{i}$.
+Since $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, we divide $\mathbf{x}^{\prime}_{i+1}$ into two parts: $(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})$ and $\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$.
+First, consider the part $(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})$.
+Substitute $\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,3})+UAT^{R}_{i}$ into it, and we get Eq.
+(7).
+Let $\mathbf{W}^{\prime}_{i+1,1}=\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}^{\prime}_{i,1}}$ and $\mathbf{b}^{\prime}_{i+1,3}=\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{b}^{\prime}_{i,1}+\mathbf{b}^{\prime}_{i+1,3}}$, we can simplify the first part to $(\mathbf{W}^{\prime}_{i+1,}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,1}UAT^{R}_{i}$.
+Since the mathematical form of $UAT^{R}_{i}$ is consistent with UAT, and we have already proven that the mathematical form of UAT remains unchanged when multiplied by a matrix, we have proven that the mathematical form of the first part is consistent with UAT.
+$(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})$ (7)
+$=$ $\{\mathbf{W}^{\prime}_{i+1,1}[(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,1})+UAT^{R}_{i}]+\mathbf{b}^{\prime}_{i+1,3}\}$
+$=$ $\{\mathbf{W}^{\prime}_{i+1,1}(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,1})+\mathbf{W}^{\prime}_{i+1,1}UAT^{R}_{i}+\mathbf{b}^{\prime}_{i+1,3}\}$
+$=$ $\{[\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{0}+\underline{(\mathbf{W}^{\prime}_{i+1,1}\mathbf{b}^{\prime}_{i,1}+\mathbf{b}^{\prime}_{i+1,3})}]+\mathbf{W}^{\prime}_{i+1,1}UAT^{R}_{i}\}$
+Next, we prove that the mathematical form of the second part, $\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, is consistent with UAT.
+By substituting $\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,3})+UAT^{R}_{i}$ into it, we obtain Eq.
+(8).
+Let $\mathbf{W}^{\prime}_{i+1,2}=\underline{\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}^{\prime}_{i,1}}$ and $\mathbf{b}^{\prime}_{i+1,2}=\underline{(\mathbf{W}^{\prime}_{i+1,2}\mathbf{b}^{\prime}_{i,1}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{W}^{\prime}_{i+1,2}UAT^{R}_{i}}$.
+Thus, the second term can be written as $\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i+1,2})$, which can be considered as a term within UAT.
+Here, $\mathbf{b}^{\prime}_{i+1,2}$ is a bias term approximated using UAT.
+$\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$ (8)
+$=$ $\mathbf{W}^{\prime}_{i+1,3}\sigma\{\mathbf{W}^{\prime}_{i+1,2}[(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,1})+UAT^{R}_{i}]+\mathbf{b}^{\prime}_{i+1,2}\}$
+$=$ $\mathbf{W}^{\prime}_{i+1,3}\sigma\{\mathbf{W}^{\prime}_{i+1,2}(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,1})+\mathbf{W}^{\prime}_{i+1,2}UAT^{R}_{i}+\mathbf{b}^{\prime}_{i+1,2}\}$
+$=$ $\mathbf{W}^{\prime}_{i+1,3}\sigma\{\underline{\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{0}+\underline{(\mathbf{W}^{\prime}_{i+1,2}\mathbf{b}^{\prime}_{i,1}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{W}^{\prime}_{i+1,2}UAT^{R}_{i}}\}$
+Since the first term of $\mathbf{x}^{\prime}_{i+1}$ has the same mathematical form as UAT, and the second term can be considered as a component within UAT, the sum of both terms still corresponds to the UAT mathematical form.
+Therefore, we have proven that the mathematical form of a multi-layer residual network is consistent with the UAT mathematical form.
+The difference lies in that some bias terms are approximated using UAT rather than being directly defined.
+C.2 Transformer to UAT
+Similarly, To illustrate relationship between multi-layer Transformer and UAT, we will first derive the general form based on Figure 2 and then compare it with the general form given in Section C.1: $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2}).$
+By examining the relationship between these forms, we can demonstrate that the mathematical form of multi-layer Transformer networks is consistent with the UAT.
+The Transformer involves two key operations: MHA and FFN.
+These operations, when expressed in matrix-vector form, correspond to Eq.(10) and Eq.(6), respectively.
+Consequently, the general term for a Transformer-based network can be written as Eq.
+(9).
+$\mathbf{x}^{\prime}_{i+1}$ $=\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i+1,3}\sigma[\mathbf{W}^{\prime}_{i+1,2}(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i})+\mathbf{b}^{\prime}_{i+1,2}]+\mathbf{b}^{\prime}_{i+1,3}$ (9)
+$=\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i+1,3}\sigma(\underline{\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}^{\prime}_{i+1,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{b}^{\prime}_{i+1,3}$
+Let $\mathbf{W}^{\prime}_{i+1,2}=\underline{\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}^{\prime}_{i+1,1}}$, so we have $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$.
+It is evident that, when compared to the general form given in Section C.1: $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, the mathematical forms are the same.
+Therefore, multi-layer Transformers are also specific implementations of the UAT.
+Similarly, to clearly present the UAT format, we use the method mentioned in Section C.1.
+Assuming the mathematical form of $\mathbf{x}^{\prime}_{i}$ is consistent with UAT, we decompose $\mathbf{x}^{\prime}_{i}$ into a primary term plus a remainder term, written as $\mathbf{x}^{\prime}_{i}=(\mathbf{W}_{i,1}\mathbf{x}_{0}+\mathbf{b}_{i,3})+UAT^{R}_{i}$.
+Substituting this into $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, we obtain Eq.
+(11).
+By setting $\mathbf{W}^{\prime}_{i+1,1}=\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}_{i,1}}$, $\mathbf{b}^{\prime}_{i+1,3}=\underline{(\mathbf{W}^{\prime}_{i+1,1}\mathbf{b}_{i,1}+\mathbf{b}^{\prime}_{i+1,3})}$, $UAT^{R}_{i}=\mathbf{W}^{\prime}_{i+1,1}UAT^{R}_{i}$, and $\mathbf{W}^{\prime}_{i+1,2}=\underline{\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}_{i,1}}$, $\mathbf{b}^{\prime}_{i+1,2}=\underline{(\mathbf{W}^{\prime}_{i+1,2}\mathbf{b}_{i,1}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{W}^{\prime}_{i+1,2}UAT^{R}_{i}}$, we get $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}_{0}+\mathbf{b}^{\prime}_{i+1,3})+UAT^{R}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}_{0}+\mathbf{b}^{\prime}_{i+1,2})$.
+Since the mathematical form of $UAT^{R}_{i}$ is consistent with multi-layer UAT, it follows that multi-layer Transformer networks are also specific implementations of UAT.
+In addition, $UAT^{R}_{i+1}=UAT^{R}_{i}+\mathbf{W}_{i+1,3}^{\prime}\sigma(\mathbf{W}_{i+1,2}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,2})$.
+So, $\mathbf{x}_{i+1}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,3})+UAT^{R}_{i+1}$, where $i=0,1,2\ldots$, $UAT_{i+1}^{R}=\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$, $\mathbf{b}^{\prime}_{j,2}=(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})+\mathbf{W}^{\prime}_{j,2}UAT_{j-1}^{R}$.
+To sum up, the $i+1$-layer Transformer can be written as:
+$\mathbf{x}_{i+1}^{\prime}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}^{\prime}+\mathbf{b}_{i+1,1})+\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$ (10)
+Other parameters will also vary with the number of layers.
+And, these changes involve only matrix multiplications or additions.
+Thus, they do not fundamentally alter the mathematical form of the UAT.
+Therefore, we will not detail the specific forms for each layer here.
+The parameters in the MHA mechanism dynamically change with the input.
+Therefore, in the corresponding UAT mathematical form for Transformers, all $\mathbf{W}^{\prime}_{i,1}$, $\mathbf{W}^{\prime}_{j,1}$ and $\mathbf{W}^{\prime}_{j,2}$ parameters for the $i$-th layer, where $j=1,...,i$, are dynamically changing with the input.
+We give some examples in Figure 3.
+$\mathbf{x}^{\prime}_{i+1}$ $=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$ (11)
+=$=\{\mathbf{W}^{\prime}_{i+1,1}[(\mathbf{W}_{i,1}\mathbf{x}_{0}+\mathbf{b}_{i,1})+UAT^{R}_{i}]+\mathbf{b}^{\prime}_{i+1,3}\}$
+$+\mathbf{W}^{\prime}_{i,3}\sigma\{\mathbf{W}^{\prime}_{i+1,2}[(\mathbf{W}_{i,1}\mathbf{x}_{0}+\mathbf{b}_{i,1})+UAT^{R}_{i}]+\mathbf{b}^{\prime}_{i+1,2}\}$
+$=[\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}_{i,1}}\mathbf{x}_{0}+\underline{(\mathbf{W}^{\prime}_{i+1,1}\mathbf{b}_{i,1}+\mathbf{b}^{\prime}_{i+1,3})}]+\mathbf{W}^{\prime}_{i+1,1}UAT^{R}_{i}]$
+$+\mathbf{W}^{\prime}_{i,3}\sigma\{\underline{\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}_{i,1}}\mathbf{x}_{0}+\underline{(\mathbf{W}^{\prime}_{i+1,2}\mathbf{b}_{i,1}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{W}^{\prime}_{i+1,2}UAT^{R}_{i}}\}$
+Figure 3: Some examples of the UAT format of multi-layer Transformer.
+The changes of parameters are represented within the dashed boxes.
+The parameters on the right of the equations indicate the original values, while those on the left represent the transformed values.
+There is no specific order of calculation for the parameters within each dashed box, but there is a top-to-bottom calculation order between different dashed boxes.
+Appendix D The Learning Process of MHA
+To more clearly illustrate the learning details of the MHA mechanism in Transformers, we present an example in Figure 4.
+First, the input consists of 10 characters, as shown in Figure 4.a.
+Then, in Figure 4.b, these characters are encoded into $n$-dimensional vectors.
+In Figure 4.c, the encoded vectors are expanded into $\mathbf{x}^{\prime}$ using Matrix-Vector methods.
+Subsequently, in Figure 4.d, we compute $\mathring{\mathbf{H}}^{\prime}=(\mathbf{W}_{HVO}^{\prime})^{T}\mathbf{x}^{\prime}$.
+Since $(\mathbf{W}_{HVO}^{\prime})^{T}$ is an almost fully dense square matrix, each element of $\mathring{\mathbf{H}}^{\prime}$ contains information from the entire input.
+This is why ICL exists in LLMs: each element can access the full context.
+While Figure 5 shows the learning process of 1D convolution.
+Compared to the holistic learning approach of MHA, 1D convolution learns by processing encoded text in chunks based on the size of the convolution kernel.
+This method is akin to learning the entire context each time but focusing on a partial set of characters of each word at a time.
+Figure 4: The learning process of MHA
+Figure 5: The learning process of 1D convolution.
+Appendix E The Comparision between Human and LLMs
+To more specifically compare the differences between humans and LLMs, we provide a simple example here.
+We prepared two questions: ”Nui zi created Nui zi dynamics” and ”Nui zi created gravitation and relativity theory.” It is important to emphasize that these sentences are entirely fictional, and there is no such person as ”Nui zi.” However, the term ”zi” has some special meanings in specific contexts that some people might know, while others may not.
+We have provided some explanations of ”zi” as interpreted by Chatgpt.4o in Figure 6.
+We have offered possible answers from different groups: a kid, a general adult, a professional researcher, and also responses from Chatgpt.4o and Qwen 2.5 to the above questions.
+Figures 7 and 8 present the answers to the two questions.
+Both humans and LLMs provide the output based on input.
+To clearly indicate which works or sentences primarily affected the generated content, we used red font to label those (this part is not generated by the model).
+Firstly, different individuals provide different answers based on their background knowledge, which can be understood as a process of generating output based on existing training parameters in the brain and the given input.
+In this process, memories about keywords lead to specific outputs, such as ”Professional researcher: 1.
+gravitation → Isaac Newton; 2.
+relativity theory → Albert Einstein.” This process can be understood as understanding the question, analyzing it, retrieving memories, and producing results.
+However, if the human brain is considered as a UAT, it generates the output based on specific inputs, such as ”gravitation → Isaac Newton” or relevant information about the theory itself.
+But the truth is that the brain never creates a ’memory warehouse’ to store that information, it just produces the output based on input.
+So the ’memory’ could be understood as something produced by input and pretrained parameters.
+Since Chatgpt.4o and Qwen 2.5 models have been trained on extensive datasets, they possess some information about ”zi,” ”Nui,” ”dynamics,” ”gravitation,” and ”relativity theory.” When these characters are inputted into the models, they use their pre-trained parameters to generate relevant outputs.
+For the first question, they provided possible inferences based on the keywords ”zi” and ”Nui” and highlighted the inconsistencies in ”Nui zi dynamics.” For the second question, which is more detailed, both Chatgpt.4o and Qwen 2.5 were able to provide more detailed contextual information (note that the models can generate additional content based on both the input and the output from the input), which can be understood as human-like understanding, analysis, and memory retrieval.
+Figure 6: The background information about ’zi’ from Chatgpt.4o.
+Figure 7: The answers about ’Nui zi created Nui zi dynamics’.
+Figure 8: The answers about ’Nui zi created gravitation and relativity theory’.

### v4

@@
-Universal Approximation Theory: The Basic Theory for Transformer-based Large Language Models
+Dynamic Universal Approximation Theory: The Basic Theory for Transformer-based Large Language Models
 Abstract
@@
 What supports the feasibility of pruning in LLMs?
-To address these key questions and explore the technical strategies within LLMs, we draw on the Universal Approximation Theory (UAT) to provide a theoretical foundation, revealing the mechanisms behind these advancements.
+To address these key questions and explore the technical strategies within LLMs, we draw on the Universal Approximation Theory (UAT) and then propose the Dynamic Universal Approximation Theory (DUAT) to provide a theoretical foundation, revealing the mechanisms behind these advancements.
 1 Introduction
@@
 Thus, endowing machines with the ability to understand and generate language marks a significant leap towards the realization of true artificial intelligence.
-The emergence of models such as the ChatGPT (Radford and Narasimhan 2018; Brown et al.
-2020; Achiam et al.
-2023), the Llama (Touvron et al.
-2023), and the PaLM (Chowdhery et al.
-2023) vividly demonstrates this point.
-A distinctive feature of LLMs is their immense parameter size (Achiam et al.
-2023; Touvron et al.
-2023; Chowdhery et al.
-2023; Chen et al.
-2021; Zeng et al.
-2022), often amounting to hundreds of billions or even trillions (for instance, GPT-3’s (Brown et al.
-2020) 175 billion parameters and PaLM’s (Chowdhery et al.
-2023) 540 billion parameters).
-This vast parameter scale lays the foundation for their exceptional language processing capabilities and enables them to exhibit almost human-like traits, such as ICL  (Brown et al.
-2020; Dong et al.
-2022), instruction following (Sanh et al.
-2021; Ouyang et al.
-2022; Wei et al.
-2021), and multi-step reasoning (Wei et al.
-2022).
-Notably, these colossal models are predominantly trained by tech giants like Google and Microsoft using large-scale GPU clusters (Zhao et al.
-2023), sparking a research fervor on how to efficiently fine-tune them with limited GPU resources.
-The advent of Lora (Hu et al.
-2021) fine-tuning technology has provided an effective pathway for this, allowing for the fine-tuning of large models under resource constraints without the need to adjust all parameters of the original model comprehensively.
-Moreover, model pruning techniques (Sun et al.
-2023; Ma, Fang, and Wang 2023) are crucial for deploying large models in resource-constrained environments, aiming to reduce the model size for operation on smaller devices.
+The emergence of models such as the ChatGPT Radford & Narasimhan (2018); Brown et al.
+(2020); Achiam et al.
+(2023), the Llama Touvron et al.
+(2023), and the PaLM Chowdhery et al.
+(2023) vividly demonstrates this point.
+A distinctive feature of LLMs is their immense parameter size Achiam et al.
+(2023); Touvron et al.
+(2023); Chowdhery et al.
+(2023); Chen et al.
+(2021); Zeng et al.
+(2022), often amounting to hundreds of billions or even trillions (for instance, GPT-3’s Brown et al.
+(2020) 175 billion parameters and PaLM’s Chowdhery et al.
+(2023) 540 billion parameters).
+This vast parameter scale lays the foundation for their exceptional language processing capabilities and enables them to exhibit almost human-like traits, such as ICL  Brown et al.
+(2020); Dong et al.
+(2022), instruction following Sanh et al.
+(2021); Ouyang et al.
+(2022); Wei et al.
+(2021), and multi-step reasoning Wei et al.
+(2022).
+Notably, these colossal models are predominantly trained by tech giants like Google and Microsoft using large-scale GPU clusters Zhao et al.
+(2023), sparking a research fervor on how to efficiently fine-tune them with limited GPU resources.
+The advent of Lora Hu et al.
+(2021) fine-tuning technology has provided an effective pathway for this, allowing for the fine-tuning of large models under resource constraints without the need to adjust all parameters of the original model comprehensively.
+Moreover, model pruning techniques Sun et al.
+(2023); Ma et al.
+(2023) are crucial for deploying large models in resource-constrained environments, aiming to reduce the model size for operation on smaller devices.
 Today, LLMs possess a diverse range of functionalities, from translation and text summarization to automatic code generation, demonstrating their versatility.
 Despite the rapid advancements of LLMs towards higher intelligence and reliability, the theoretical foundations behind these models remain largely unexplored and shrouded in uncertainty.
-The scientific community is actively seeking to uncover the intrinsic mechanisms behind their powerful capabilities, including the analysis of ICL mechanisms (Xie et al.
-2021; Min et al.
-2022).
-In this context, we establish the UAT (Cybenko 2007; Hornik, Stinchcombe, and White 1989) as the mathematical essence of Transformer-based LLMs and use the perspective of UAT to elucidate key technologies and phenomena within LLMs, aiming to provide explanations from a new angle.
-While previous studies have explored the relationship between Transformers and UAT (Yun et al.
-2019, 2020; Kratsios et al.
-2021; Alberti et al.
-2023), these studies are based on a series of complex assumptions and derivations, lacking generality and failing to intuitively address some of the issues present in LLMs.
-In contrast, this paper aims to provide proof that bridges Transformers and UAT, grounded in the foundational mathematical form of UAT as established by Cybenko; Hornik, Stinchcombe, and White.
-Our proof is characterized by its simplicity, absence of assumptions, and generality, making it applicable to Convolutional Neural Networks (CNNs) as well (Wang and Li 2024b).
-Our approach not only provides intuitive answers to some of the challenges within LLMs but also aids in the design of deep learning networks from a theoretical standpoint (Wang and Li 2024a).
-These capabilities are not addressed by previous theories.
+The scientific community is actively seeking to uncover the intrinsic mechanisms behind their powerful capabilities, including the analysis of ICL mechanisms Xie et al.
+(2021); Min et al.
+(2022).
+In this context, we leverage UAT and then establish the DUAT as the mathematical essence of Transformer-based LLMs and use the perspective of DUAT to elucidate key technologies and phenomena within LLMs, aiming to provide explanations from a new angle.
+Although previous studies have explored the relationship between Transformer Vaswani2017AttentionIA and the UAT Yun et al.
+(2019; 2020); Kratsios et al.
+(2021); Alberti et al.
+(2023), they rely on a series of complex assumptions and derivations, which limits their generalizability and fails to address certain issues in LLMs in a straightforward manner.
+In contrast, this paper directly demonstrates the connection between multi-layer Transformer and UAT by representing it in a mathematical format.
+Cybenko (1989) and Hornik et al.
+Hornik et al.
+(1989) established the fundamental mathematical form of UAT, which we extend to the DUAT by leveraging the residual structure of Transformers.
+This form can account for several key issues in LLMs, such as why LLMs are capable of handling both translation and generation tasks simultaneously, and whether LLMs possess memory capacity (Wang & Li, 2024a).
+Moreover, DUAT is equally applicable to convolutional neural networks (CNNs) Wang & Li (2024b).
+Our approach not only provides intuitive explanations for some of the challenges within LLMs but also offers theoretical insights that aid in the design of deep learning networks Wang & Li (2024c), covering capabilities that previous theories have not addressed.
 Our contributions are as follows:
-  • We prove that the Transformer is the tangible embodiment of UAT.
-  • We deliver a rigorous scientific explanation of Transformer-based LLMs through the UAT.
+  • We further develop the UAT to DUAT.
+  • We prove that the Transformer is the tangible embodiment of DUAT.
+  • We deliver a rigorous scientific explanation of Transformer-based LLMs through the DUAT.
   • We explain the characteristics of LLMs, such as ICL, instruction following, multi-step reasoning, and the technologies applied within LLMs like Lora, pruning, and LLMs’ strong generalization capabilities.
-Our article is structured as follows: In Section 2, we begin by introducing the UAT and propose that to demonstrate the Transformer’s adherence to UAT, it is necessary to show that both Linear and Multi-Head Attention (MHA) can be represented in the form of matrix-vector multiplication.
-In Section 3, we establish that the Transformer belongs to the category of UAT.
-Our proof strategy unfolds starting with Section 3.1, where we present the idea to express Linear and MHA as matrix-vector products.
-In Section 3.2, we detail the computational methods required for the proof.
-Sections 3.3 and 3.4 are dedicated to proving that both Linear and MHA can indeed be represented as matrix-vector operations.
-Further in Section 4, we leverage UAT to theoretically elucidate some fundamental issues (Generalization 4.1, ICL 4.2) and techniques (Pruning 4.3 and LoRA 4.4) associated with LLMs.
-Finally, we conclude with a summary of the existing challenges and potential future developments for LLMs in Section 4.5.
-2 The Universal Approximation Theory
-Up to now, the UAT (Cybenko 2007) is the most widely recognized fundamental theory in deep learning.
-However, this theorem is only applicable to the simplest form of neural networks, the multilayer perceptron (MLP) (Cybenko 2007; Popescu et al.
-2009).
-Due to the increased complexity of Transformer networks, they cannot be mathematically expressed in the same form as UAT.
-Thus, the theory has not yet been extended to Transformer networks.
-Our aim in this paper is to unify Transformer networks under the framework of UAT, thereby standardising their mathematical representation.
-Before unifying their mathematical forms, we provide a brief overview of the theorem, which was initially proposed by Cybenko (2007).
-This theorem encompasses numerous conclusions and proof details.
-Although it has been further developed, its fundamental mathematical form remains unchanged.
-Therefore, this paper explains the theory based on the UAT form presented by Cybenko (2007).
-Theorem 2 from Cybenko (2007) states that if $\sigma$ is any continuous sigmoidal function, then finite sums of the following form:
-$G(\mathbf{x})=\sum_{j=1}^{N}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$ (1)
+The structure of this paper is as follows: in Section 2.1, we first introduce the UAT and propose that to prove Transformers adhere to the DUAT framework, it is necessary to demonstrate that both the linear layer and the multi-head attention mechanism (MHA) can be expressed in the form of matrix-vector multiplication.
+Then, in Section 2.2, we present a general framework for representing various operations in deep learning as matrix-vector multiplication.
+In Section 3, we establish that Transformers fall within the scope of DUAT.
+Our proof strategy involves representing the linear layer and MHA as matrix-vector operations.
+Sections 3.1 and 3.2 provide detailed proofs that the linear layer and MHA can indeed be expressed in this form.
+In Section 3.3, we present the DUAT representation of a multi-layer Transformer network.
+In Section 4, we leverage the DUAT framework to theoretically elucidate several fundamental questions (e.g., generalization ability, 4.1, and in-context learning capability, 4.2) and technical aspects (e.g., feasibility of pruning, 4.3, and the effectiveness of the LoRA scheme, 4.4) related to large language models (LLMs).
+Finally, in Section 6, we summarize the current challenges of LLMs and explore potential directions for future research.
+2 The Universal Approximation Theory and Matrix-Vector Method
+2.1 The Universal Approximation Theory
+The Universal Approximation Theorem (UAT) (Cybenko, 1989) remains one of the most widely recognized foundational theories in the field of deep learning to date.
+However, the theorem primarily applies to the simplest forms of neural networks—single-layer or multilayer perceptrons (MLPs) (Cybenko, 1989; Popescu et al., 2009).
+Due to the significantly increased complexity of Transformer networks, they cannot be expressed mathematically in the same manner as those governed by UAT.
+Consequently, the extension of UAT to Transformer networks has yet to be established.
+This paper aims to develop the DUAT (an advance of UAT) and explore Transformer networks within this framework to standardize their mathematical representation.
+Before achieving this, we will briefly revisit the UAT as initially proposed by Cybenko (1989).
+The theorem encompasses a wealth of conclusions and detailed proofs.
+Although it has been further developed over time, its core mathematical form remains unchanged.
+Thus, this paper interprets the theory based on the UAT form articulated by Cybenko (1989).
+According to Theorem 2 in Cybenko (1989), if $\sigma$ is any continuous sigmoid function, then a finite sum of the following form:
+$G(\mathbf{x})=\sum_{j=1}^{N}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+b_{j}\right)$ (1)
 is dense in $C\left(\mathbf{I}_{n}\right)$.
-Here, $\mathbf{W}_{j}\in\mathbb{R}^{n}$ and $\alpha_{j},\theta\in\mathbb{R}$ are fixed.
+Here,$\mathbf{x}\in\mathbf{I}_{n}$, $\mathbf{W}_{j}\in\mathbb{R}^{n}$ and $\alpha_{j},b\in\mathbb{R}$ are fixed.
 For any $f\in C\left(\mathbf{I}_{n}\right)$ and $\varepsilon>0$, there exists a function $G(\mathbf{x})$:
-$|G(\mathbf{x})-f(\mathbf{x})|<\varepsilon\quad\text{ for all }\quad\mathbf{x}\in\mathbf{I}_{n}.$ (2)
+$|G(\mathbf{x})-f(\mathbf{x})|<\varepsilon$ (2)
 This implies that, when $N$ is sufficiently large, a neural network can approximate any continuous function on a closed interval.
-Hornik, Stinchcombe, and White (1989) further demonstrates that multilayer feedforward networks also conform to the UAT, capable of approximating arbitrary Borel measurable functions.
-Observing Equation (1), where the function $G(\mathbf{x})$ yields a scalar output in $\mathbb{R}$, the scenario expands naturally when $G(\mathbf{x})$ maps to $\mathbb{R}^{m}$, requiring the approximation in each dimension.
-It becomes evident that to accommodate this multidimensional output, a simple extension to Equation (1) suffices: the transformation matrix $\mathbf{W}_{j}$ is revised to the space $\mathbb{R}^{n\times m}$, the bias term $\theta_{j}$ is recast as a vector in $\mathbb{R}^{m}$, and $\alpha_{j}$ is reshaped into a matrix.
+Hornik et al.
+(1989) further demonstrates that multilayer feedforward networks also conform to the UAT, capable of approximating arbitrary Borel measurable functions.
+Observing Eq.
+1, where the function $G(\mathbf{x})$ yields a scalar output in $\mathbb{R}$, the scenario expands naturally when $G(\mathbf{x})$ maps to $\mathbb{R}^{m}$, requiring the approximation in each dimension.
+It becomes evident that to accommodate this multidimensional output, a simple extension to Eq.
+1 suffices: the transformation matrix $\mathbf{W}_{j}$ is revised to the space $\mathbb{R}^{n\times m}$, the bias term $b_{j}$ is recast as a vector in $\mathbb{R}^{m}$, and $\alpha_{j}$ is reshaped into a matrix.
 Nevertheless, in these formulas, the theorem does not straightforwardly apply to Transformer architectures.
-However, if we succeed in reformulating both the Linear ($\mathbf{x}|\mathbf{W}$) and MHA ($\mathbf{x}|\mathbf{W}$) components into a unified representation, $\mathbf{W}^{\prime}\mathbf{x}^{\prime}$, where $\mathbf{W}$ and $\mathbf{x}$ represent the parameters and inputs for each component respectively, and $\mathbf{W}^{\prime}$ is derived from $\mathbf{W}$ while $\mathbf{x}^{\prime}$ signifies a column vector derived from $\mathbf{x}$ (for details, refer to Section 3.1).
-It is easy to prove that multi-layer Transformer networks are concrete embodiments of UAT.
-Consequently, our objective is to represent Linear and MHA architectures as a matrix-vector multiplication similar to Eq.
-(3).
-This approach offers two key benefits: Firstly, it allows for a fundamental mathematical investigation of deep learning models.
-Secondly, it provides a more intuitive and objective basis for comparing differences among various models.
-On this foundation, we propose to transform Linear and MHA into the matrix-vector format.
-3 UAT for Transformer
-In the previous section, we have outlined our fundamental goal to unify the Transformer model under the framework of UAT by converting Linear and MHA operations in the Transformer into matrix-vector forms.
-In this section, we will use the Matrix-Vector method to transform Transformer operations into matrix times vector forms.
-3.1 Matrix-Vector Method
+However, if we succeed in reformulating both the Linear ($\mathbf{x}|\mathbf{W}$) and MHA ($\mathbf{x}|\mathbf{W}$) components into a unified representation, $\mathbf{W}^{\prime}\mathbf{x}^{\prime}$, where $\mathbf{W}$ and $\mathbf{x}$ represent the parameters and inputs for each component respectively, and $\mathbf{W}^{\prime}$ is derived from $\mathbf{W}$ while $\mathbf{x}^{\prime}$ signifies a column vector derived from $\mathbf{x}$.
+It is easy to build the connection between multi-layer Transformer networks and UAT.
+2.2 The Matrix-Vector Method
 Figure 1: The transformation process of the Matrix-Vector Method.
@@
 This method constitutes a strategic realignment of input data and corresponding parameters of various transformations within the network, as illustrated in Figure 1.
-The underlying principle is as follows: both the input ($\mathbf{x}$) and output ($\mathbf{y}$) data are reconfigured uniformly through a transformation $\mathbf{T_{D}}$ into column vectors ($\mathbf{x}^{\prime}$ and $\mathbf{y}^{\prime}$, while parameter tensor ($\mathbf{W}$) is reorganized into matrix form ($\mathbf{W}^{\prime}$).
-A critical requirement is that identical operations throughout the LLMs network adhere to the same restructuring scheme, thereby eliminating the need for additional transformations on intermediate feature data; they can directly be represented as column vectors.
+The underlying principle is as follows: both the input ($\mathbf{x}$) and output ($\mathbf{y}$) data are reconfigured uniformly through a transformation $\mathbf{T_{D}}$ into column vectors ($\mathbf{x}^{\prime}$ and $\mathbf{y}^{\prime}$, while parameter tensor $\mathbf{W}$ is reorganized into matrix form $\mathbf{W}^{\prime}$ by $\mathbf{T_{P}}$.
+$\mathbf{T_{D}}$ and $\mathbf{T_{P}}$ are transformations and their purpose is to reorganize the input $\mathbf{x}^{\prime}$ and $\mathbf{y}^{\prime}$ into column vectors and $\mathbf{W}^{\prime}$ fulfill $\mathbf{W}^{\prime}\mathbf{x}^{\prime}=\mathbf{y}^{\prime}$.
+A critical requirement is that the input and output throughout the LLMs network adhere to the same restructuring scheme, thereby eliminating the need for additional transformations on intermediate feature data; they can directly be represented as column vectors.
 Additionally, by default, matrix variables in the original formulas are represented in bold, such as $\mathbf{x}$, and in the Matrix-Vector form, corresponding variables are denoted with a prime symbol () in the upper right corner, such as $\mathbf{x}^{\prime}$.
@@
 The Matrix-Vector Method can be succinctly encapsulated as follows:
-$\mathbf{y}=T(\mathbf{x}|\mathbf{W})\rightarrow\mathbf{y}^{\prime}=\mathbf{W}^{\prime}\mathbf{x}^{\prime}$ (3)
+$\mathbf{y}=T(\mathbf{x}|\mathbf{W})\mapsto\mathbf{y}^{\prime}=\mathbf{W}^{\prime}\mathbf{x}^{\prime}$ (3)
+where $T$ represents a general transformation, like MHA.
 It is obvious that $T_{D}$ and $T_{P}$ are not fixed, we could design various kinds of ways to do those.
 For convenience, we present a methodology tailored for Linear in Section 3.1 and an approach for MHA in Section 3.2, thereby illustrating the adaptability and application of the Matrix-Vector Method across different components of the Transformer architecture.
-3.2 Diamond Matrix Multiplication
-While our goal is to transform Linear and MHA modules into a matrix-vector format, this process is complex.
-In order to describe this transformation process clearly, we propose a new computational method called the Diamond Multiplication Method, denoted by the symbol $\diamond$.
-(The purpose of proposing this method is to clearly illustrate the calculation process in turning MHA into Matrix-Vector format.
-If we use direct matrix multiplication, some computational details cannot be visually distinguished.)
-$\left(\begin{array}[]{cccc}w_{1,1}&w_{1,2}&\cdots&w_{1,n}\\ w_{2,1}&w_{2,2}&&w_{2,n}\\ \vdots&\vdots&\vdots&\vdots\\ w_{m,1}&w_{m,2}&&w_{m,n}\end{array}\right)\diamond\left(\begin{array}[]{c}x_{1}\\ x_{2}\\ \vdots\\ x_{m}\end{array}\right)$ (4)
-$=\left(\begin{array}[]{c}w_{1}x_{1}+w_{21}x_{2}+\cdots w_{m,1}x_{m}\\ w_{1,2}x_{1}+w_{2,2}x_{2}+\cdots w_{m,2}x_{m}\\ \vdots\\ w_{1,n}x_{1}+w_{2,n}x_{2}+\cdots w_{mn}x_{m}\end{array}\right)$
-Let $\mathbf{W}\in\mathbb{R}^{(m,n)}$ and $\mathbf{x}\in\mathbb{R}^{(m,1)}$, the Diamond Matrix Multiplication is defined as $\mathbf{y}=\mathbf{W}\diamond\mathbf{x}$, where $\mathbf{y}\in\mathbb{R}^{(m,1)}$.
-The computation procedure involves element-wise multiplication and summation of corresponding elements from left to right across the columns of matrix $\mathbf{W}$ with vector $\mathbf{x}$.
-The sum of the multiplication of the $i$th column of $\mathbf{W}$ with $\mathbf{x}$ is assigned as the $i$th element of $\mathbf{y}$.
-A detailed calculation process is provided in Eq.
-(4).
-The computation process for $\mathbf{y}=\mathbf{x}\diamond\mathbf{W}$ is identical to that of $\mathbf{W}\diamond\mathbf{x}$, implying that $\mathbf{W}\diamond\mathbf{x}=\mathbf{x}\diamond\mathbf{W}$.
-Furthermore, based on its computation method, it can be inferred that the Diamond Matrix Multiplication is related to conventional matrix multiplication, specifically: $\mathbf{W}\diamond\mathbf{x}=\mathbf{W}^{T}\mathbf{x}$.
-Additionally, when $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are square matrices, we can derive the following relationships: $\mathbf{W}_{1}\diamond[\mathbf{x}\diamond\mathbf{W}_{2}]=\mathbf{W}_{2}^{T}\diamond\mathbf{W}_{1}\diamond\mathbf{x}$ and $[\mathbf{W}_{1}\diamond\mathbf{x}]\diamond\mathbf{W}_{2}=\mathbf{W}_{1}^{T}\diamond\mathbf{W}_{2}\diamond\mathbf{x}$.
-Detailed derivations can be found in Appendix A.
+Figure 2: This diagram shows the differences between an UAT and a DUAT.
+2.3 The introduction of DUAT
+Although the UAT function has strong approximation capabilities, it has a fatal flaw: it can only fit one function within a closed interval.
+This limitation results in poor generalization.Therefore, to enable the model to fit different functions within a closed interval based on the input, we need to grant the parameters in the UAT the ability to dynamically adjust according to the input, as defined in Eq.
+4.
+$\mathbf{y}_{i}={G(\mathbf{x}_{i}|\mathbf{W(\mathbf{x}_{i})},\mathbf{b(\mathbf{x}_{i})})}$ (4)
+where $\mathbf{W}=\mathbf{W}_{1}\cdots\mathbf{W}_{N}$ and $\mathbf{b}=b_{1}\cdots b_{N}$igure 2 illustrates the comparison between UAT and DUAT.
+On the left side, Figure 2.a demonstrates the deficiency in data recognition.
+The input to the network is shown, with only the initial part differing slightly.
+This can be likened to tasks in LLMs, such as translating, continuing, or summarizing the same passage, where only the initial text segment differs while the rest remains the same.
+Figure 2.b provides a schematic illustration of the output for the input shown in Figure 2.a.
+Because the UAT can only fit a single function, the output closely follows the trend of the input, resulting in nearly identical outcomes regardless of the slight differences in the initial part.
+Essentially, the output remains the same regardless of the prompt.
+In contrast, Figure 2.c shows the schematic output of the DUAT (Dynamic Universal Approximation Theorem) function.
+In DUAT, some or all parameters are functions of the input, allowing it to produce different outputs based on different prompts.
+This flexibility enables DUAT to generate varied results.
+On the right side, Figure 2 compares UAT and DUAT through the distribution of data.
+Since the parameters of UAT are fixed, it can only fit a fixed data distribution.
+Given that we typically train with large datasets, the goal for UAT is to fit as much data as possible, which often results in fitting only the densest part of the data distribution, represented in purple.
+On the other hand, DUAT can dynamically adjust its parameters based on the input, allowing it to fit multiple different distributions simultaneously.
+This comparison highlights the advantages of DUAT in handling more complex and diverse data scenarios, making it a more versatile and powerful tool for various applications.
+3 DUAT for Transformer
+In this section, we will apply the matrix-vector approach to convert the operations in the Transformer Vaswani2017AttentionIA into matrix-vector multiplication form, and then explain the defination of DUAT and demonstrate that the Transformer is essentially a DUAT function.
 3.1 Matrix-Vector Method for Linear
@@
 Figure 3 illustrates this process: Figure 3.a shows the linear transformation of multi-channel input: $\mathbf{W}\mathbf{x}=\mathbf{y}$.
-Figure 2.b provides a specific example of Figure 2.a, Figure 2.c converts the linear transformation in Figure 2.b into the corresponding matrix-vector representation: $\mathbf{W}^{\prime}\diamond\mathbf{x}^{\prime}=\mathbf{y}^{\prime}$.
+Figure 3.b provides a specific example of Figure 3.a, Figure 3.c converts the linear transformation in Figure 3.b into the corresponding matrix-vector representation: $\mathbf{W}^{\prime}\mathbf{x}^{\prime}=\mathbf{y}^{\prime}$.
 Thus, the linear transformation can be represented in matrix-vector form as follows:
-$\mathbf{x}_{i+1}=\mathbf{W}_{i}\mathbf{x}^{i}\rightarrow\mathbf{x}_{i+1}^{\prime}$ $=\mathbf{W}_{i}^{\prime}\diamond\mathbf{x}_{i}^{\prime}=(\mathbf{W}_{i}^{\prime})^{T}\mathbf{x}_{i}^{\prime}$ (5)
+$\mathbf{x}_{i+1}=\mathbf{W}_{i}\mathbf{x}_{i}\mapsto\mathbf{x}_{i+1}^{\prime}=\mathbf{W}^{\prime}_{i}\mathbf{x}^{\prime}_{i}$ (5)
 Here, $\mathbf{x}_{i}\in\mathbb{R}^{(N,M)}$ and $\mathbf{x}_{i+1}\in\mathbb{R}^{(N,M)}$ represent the input and output of layer $i$, respectively, while $\mathbf{W}_{i}\in\mathbb{R}^{(N,N)}$ represents the parameters of layer $i$.
@@
 $FFN(\mathbf{x})$ $=\mathbf{W}_{2}\sigma(\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1})+\mathbf{b}_{2}$ (6)
-$\rightarrow\mathbf{W}_{2}^{\prime}\sigma(\mathbf{W}_{1}^{\prime}\mathbf{x}^{\prime}+\mathbf{b}_{1}^{\prime})+\mathbf{b}_{2}^{\prime}$
-Figure 3: The Transformer process.
-Figure 4: The process of transforming $\hat{\mathbf{H}}=\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}}_{8})$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{HV}^{\prime}\diamond\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$.
-Figure 5: This diagram illustrates the process of transforming $\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}})\mathbf{W}_{O}$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{HVO}^{\prime}\diamond\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, different colors in the diagram correspond to different variables.
+$\mapsto\mathbf{W}_{2}^{\prime}\sigma(\mathbf{W}_{1}^{\prime}\mathbf{x}^{\prime}+\mathbf{b}_{1}^{\prime})+\mathbf{b}_{2}^{\prime}$
 3.2 Matrix-Vector Method for MHA
@@
 $=$ $\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}=\hat{\mathbf{H}}\mathbf{W}_{O}$
-$\operatorname{Attention}(\mathbf{x}_{i}\mathbf{W}_{Qi}=\mathbf{Q}_{i},\mathbf{x}_{i}\mathbf{W}_{Ki}=\mathbf{K}_{i},$ (8)
-$\mathbf{x}_{i}\mathbf{W}_{Vi}=\mathbf{V}_{i})=\operatorname{softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{T}}{\sqrt{M}}\right)\mathbf{V}_{i}$
+$\operatorname{Attention}(\mathbf{x}_{i}\mathbf{W}_{Qi}=\mathbf{Q}_{i},\mathbf{x}_{i}\mathbf{W}_{Ki}=\mathbf{K}_{i},\mathbf{x}_{i}\mathbf{W}_{Vi}=\mathbf{V}_{i})$ (8)
+$=\operatorname{softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{T}}{\sqrt{M}}\right)\mathbf{V}_{i}$
 $=\mathbf{H}_{i}\mathbf{V}_{i}=\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]=\hat{\mathbf{H}}_{i}$
@@
 Next, we give the proof of this process.
+Figure 4: The Transformer process.
 Our objective is to express the MHA as $(\mathbf{W}^{\prime}_{HVO})^{T}\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$.
@@
 In Eq.
-(7), $\hat{\mathbf{H}}=\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)$ describes an engineering process that requires mathematical representation.
+7, $\hat{\mathbf{H}}=\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)$ describes an engineering process that requires mathematical representation.
 The learning process for the input $\mathbf{x}$ primarily consists of two parts: $\mathbf{V}_{1},\ldots,\mathbf{V}_{h}$ and $\mathbf{H}_{1},\ldots,\mathbf{H}_{h}$.
@@
 Figure 5.b depicts $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}],i=1,2$.
-In Figure 4.c we convert Figure 4.b into the matrix-vector form $\mathbf{W}_{HV}^{\prime}\diamond\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{HV}^{\prime}$ is generated from $\mathbf{H}_{1},\mathbf{H}_{2}$ and $\mathbf{W}_{V1},\mathbf{W}_{V2}$.
+In Figure 5.c we convert Figure 5.b into the matrix-vector form $(\mathbf{W}_{HV}^{\prime})^{T}\mathbf{x}^{\prime}=(\hat{\mathbf{H}}^{\prime})^{T}$, where $\mathbf{W}_{HV}^{\prime}$ is generated from $\mathbf{H}_{1},\mathbf{H}_{2}$ and $\mathbf{W}_{V1},\mathbf{W}_{V2}$.
 More details can be found in Appendix B.
+Figure 5: The process of transforming $\hat{\mathbf{H}}=\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}}_{8})$ in the MHA into its corresponding matrix-vector form $(\mathbf{W}_{HV}^{\prime})^{T}\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$.
 Figure 6 illustrates the parameter transformation scenario after incorporating $\mathbf{W}_{O}$.
-Figure 5.a demonstrates an example of $\mathring{\mathbf{H}}=\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}$, while Figure 5.b rewrites Figure 5.a as $\hat{\mathbf{H}}^{\prime}\diamond\mathbf{W}_{O}^{\prime}=\mathring{\mathbf{H}}^{\prime}$.
-Consequently, based on Figures 4 and 5, the entire MHA can be expressed as $\mathbf{W}_{HVO}^{\prime}\diamond\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{HVO}^{\prime}=\mathbf{W}_{HV}^{T}\diamond\mathbf{W}_{O}^{\prime}$ (Based on Figures 5, we can get $\mathbf{W}_{HVO}^{\prime}$, but the resulting expression is too complex and not easy to understand, so we did not compute the final result.
-However, it is easy to analyze that as long as the variables in $\mathbf{W}_{HV}^{T}$ and $\mathbf{W}_{O}^{\prime}$ are non-zero, the result is essentially a dense matrix).
+Figure 6.a demonstrates an example of $\mathring{\mathbf{H}}=\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}$, while Figure 6.b rewrites Figure 6.a as $(\mathbf{W}_{O}^{\prime})^{T}\hat{\mathbf{H}}^{\prime}=\mathring{\mathbf{H}}^{\prime}$.
+Consequently, based on Figures 5 and 6, the entire MHA can be expressed as $(\mathbf{W}_{HVO}^{\prime})^{T}\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{HVO}^{\prime}=\mathbf{W}_{O}^{\prime}\mathbf{W}_{HV}^{\prime}$ (Based on Figures 6, we can get $\mathbf{W}_{HVO}^{\prime}$, but the resulting expression is too complex and not easy to understand, so we did not compute the final result.
+However, it is easy to analyze that as long as the variables in $\mathbf{W}_{HV}$ and $\mathbf{W}_{O}^{\prime}$ are non-zero, the result is essentially a dense matrix).
 So the whole MHA can be written as:
@@
 $\mathring{\mathbf{H}}^{\prime}=\mathbf{W}^{\prime}\mathbf{x}^{\prime}$ (10)
-3.5 The UAT Format of Multi-Layer Transformer
+Figure 6: This diagram illustrates the process of transforming $\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}})\mathbf{W}_{O}$ in the MHA into its corresponding matrix-vector form $(\mathbf{W}_{HVO}^{\prime})^{T}\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, different colors in the diagram correspond to different variables.
+3.3 The DUAT Format of Multi-Layer Transformer
 The Transformer architecture is founded on two pivotal components: FFN and MHA.
-In Sections 3.3 and 3.4, we have showcased the matrix-vector representations for both FFN and MHA.
-In this section, we delve into why multi-layer Transformer is the implementation of UAT.
-Based on Equations 6 and 10, we can demonstrate that $i+1$-th layer ($i=0,1,2,\ldots$) Transformer is a concrete realization of the UAT, which can be written as (see Appendix C for more details):
-$\mathbf{x}_{i+1}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,1})+\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$ (11)
-where $\mathbf{x}_{i+1}$ is the output of ${i+1}$-th layer, $\mathbf{x}_{0}$ is the input of the network, $\mathbf{b}^{\prime}_{j,2}=(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})+\mathbf{W}^{\prime}_{j,2}UAT_{j-1}^{R}$, where $UAT_{j-1}^{R}=\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,3}\sigma(\mathbf{W}^{\prime}_{k,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,2})$.
-The term $\mathbf{b}^{\prime}_{j,2}$ is approximated by the $j$ layer of UAT with $\mathbf{x}_{0}$ as input.
-This enhances the model’s ability to dynamically adjust functions based on input.
+In Sections 3.1 and 3.2, we have showcased the matrix-vector representations for both FFN and MHA.
+In this section, we delve into why multi-layer Transformer is the implementation of DUAT.
+Based on Eq 6 and 28, we can demonstrate that $i+1$-th layer ($i=0,1,2,\ldots$) Transformer can be written as (see Appendix C for more details):
+$\mathbf{x}_{i+1}=$ $(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,1})$ (11)
+$+\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$
+where $i=0$, the parameters are defined as follows: $\mathbf{W}^{\prime}_{1,1}=\mathbf{W}^{\prime}_{1,1}$, $\mathbf{b}^{\prime}_{1,3}=\mathbf{b}^{\prime}_{1,3}$, $\mathbf{W}^{\prime}_{1,3}=\mathbf{W}^{\prime}_{1,3}$, $\mathbf{W}^{\prime}_{1,2}=\mathbf{W}^{\prime}_{1,2}\mathbf{W}^{\prime}_{1,1}$, and $\mathbf{b}^{\prime}_{1,2}=\mathbf{b}^{\prime}_{1,2}$.
+We define the parameters for cases where $i\geq 1$.
+For each $j=1,2,\ldots,i$, the updates are as follows: $\mathbf{W}^{\prime}_{j+1,1}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{W}_{j,1}$, $\mathbf{b}^{\prime}_{j+1,3}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{b}_{j,3}+\mathbf{b}^{\prime}_{j+1,3}$, $\mathbf{W}^{\prime}_{j+1,2}=\mathbf{W}^{\prime}_{j+1,2}\mathbf{W}_{j,1}$, and $\mathbf{W}^{\prime}_{j,3}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{W}^{\prime}_{j,3}$.
+Additionally, for $j=2,\ldots,i+1$, the bias terms $\mathbf{b}^{\prime}_{j,2}$ are updated according to:
+$\mathbf{b}^{\prime}_{j,2}=$ $(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})$ (12)
+$+$ $\mathbf{W}^{\prime}_{j,2}\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,3}\sigma(\mathbf{W}^{\prime}_{k,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,2}).$
+The term $\mathbf{b}^{\prime}_{j,2}$ is approximated by the $j$ layer UAT with $\mathbf{x}_{0}$ as input.
 In the MHA mechanism, the parameters change dynamically with the input.
-Therefore, in the formula above, all $\mathbf{W}^{\prime}_{j,1}$, $\mathbf{W}^{\prime}_{j,2}$, and $\mathbf{W}^{\prime}_{j,3}$ parameters in layer $i$, where $j=1,\ldots,i$, are dynamically adjusted based on the input.
+Therefore, in the formula above, all $\mathbf{W}^{\prime}_{j,1}$ and $\mathbf{W}^{\prime}_{j,2}$ for $i+1\geq j\geq 1$, and $\mathbf{W}^{\prime}_{j,3}$ for $i+1>j>1$ in layer $i$, are dynamically adjusted based on the input.
+According to Eq.
+11, we know the mathematical format of a multi-layer Transformer is the same as UAT, but the parameters in the multi-layer Transformer’s UAT are not fixed, they will change with input.
+So we call the UAT format of parameters dynamically adjust based on input as DUAT.
 4 Discussion
-Leveraging on our proof in the previous section that the Transformer is the tangible embodiment of UAT, in this section we address the following critical problems and explore the technical strategies of LLMs: Why does the Transformer architecture possess such power in enabling intelligent language models, such as translation and programming?
+Leveraging on our proof in the previous section that the Transformer is the tangible embodiment of DUAT, in this section we address the following critical problems and explore the technical strategies of LLMs: Why does the Transformer architecture possess such power in enabling intelligent language models, such as translation and programming?
 What enables LLMs’ capacity for ICL?
@@
 4.1 What Makes Transformer So Powerful in LLMs?
-Theoretically, we have established that Transformer networks are concrete implementations of the UAT, enabling them to approximate any Borel measurable function.
+Theoretically, we have established that Transformer networks are concrete implementations of the DUAT, enabling them more powerful approximation ability than UAT.
 While UAT provides powerful function approximation capabilities, it inherently lacks the ability to approximate multiple functions simultaneously.
@@
 The MHA and residual mechanisms in Transformers equip LLMs with the ability to dynamically approximate relevant functions according to the input.
-Specifically, MHA allows for the dynamic adjustment of the weight parameters in UAT in response to the input, while the residual mechanism supports the dynamic approximation of bias terms, as illustrated in Figure 6.
+Specifically, MHA allows for the dynamic adjustment of the weight parameters in UAT in response to the input, while the residual mechanism supports the dynamic approximation of bias terms.
 This theoretical foundation enables Transformer-based LLMs to handle a wide range of tasks, including translation, continuation, summarization, code generation, and solving mathematical problems.
-Figure 6: This diagram vividly demonstrates the output differences between an MLP and a Transformer when given the same input content (represented by the blue line in a) with only slight variations (represented by the black line in a).
-Since the input content is nearly identical and the MLP’s parameters are fixed, the MLP tends to produce the same output (b), influenced by the overall input function trend.
-However, the Transformer, with its ability to dynamically adapt to the input, generates different outputs based on the input variations.
-In the context of LLMs, the blue line in a can be understood as a consistent piece of text, while the black line represents different prompts, such as translation, continuation, or summarization.
 Furthermore, the MHA mechanism can capture global information (as shown in Figure 6, where each element in the output $\mathring{\mathbf{H}}^{\prime}$ contains global information), which helps generate content that is consistent with the context.
@@
 So, how does this contextual interaction capability arise within LLMs?
-The formula $\mathring{\mathbf{H}}=\mathbf{W}^{\prime}\mathbf{x}^{\prime}$ in Figure 5 reveals this mode of contextual interaction.
-Since $\mathbf{W}^{\prime}$ represents a dense matrix (almost devoid of zero elements and whose internal elements are highly correlated), each element in $\mathring{\mathbf{H}}$ encapsulates comprehensive information from both preceding and subsequent contexts.
+The formula $\mathring{\mathbf{H}}=(\mathbf{W}^{\prime}_{HVO})^{T}\mathbf{x}^{\prime}$ in Figure 6 reveals this mode of contextual interaction.
+Since $\mathbf{W}^{\prime}_{HVO}$ represents a dense matrix (almost devoid of zero elements and whose internal elements are highly correlated), each element in $\mathring{\mathbf{H}}$ encapsulates comprehensive information from both preceding and subsequent contexts.
 This learning of holistic contextual information constitutes the foundation of contextual interaction within LLMs.
@@
 To understand this, we can directly analyze it from the perspective of the formula underlying the UAT:
-$|\sum_{j=1}^{N}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)-f(\mathbf{x})|<\varepsilon\quad.$ (12)
+$|\sum_{j=1}^{N}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+b_{j}\right)-f(\mathbf{x})|<\varepsilon\quad.$ (13)
 for all $\mathbf{x}\in\mathbf{I}_{n}$.
-Let’s assume $\Lambda=\{1,2...N\},\Lambda_{1}\cup\Lambda_{2}=\Lambda,\Lambda_{1}\cap\Lambda_{2}=\emptyset$ and $|\sum_{j\in\Lambda_{1}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)|\to 0$.
+Let’s assume $\Lambda=\{1,2...N\},\Lambda_{1}\cup\Lambda_{2}=\Lambda,\Lambda_{1}\cap\Lambda_{2}=\emptyset$ and $|\sum_{j\in\Lambda_{2}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)|\to 0$.
 Then we have:
-$|\sum_{j\in\Lambda_{1}}\alpha_{j}\sigma\left(\mathbf{W}j^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$ (13)
-$+$ $\sum_{j\in\Lambda_{2}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)-f(\mathbf{x})|<\varepsilon$
-Since $|\sum_{j\in\Lambda_{1}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)|\to 0$, we have the following inequality:
+$|\sum_{j\in\Lambda_{1}}\alpha_{j}\sigma\left(\mathbf{W}j^{\mathrm{T}}\mathbf{x}+b_{j}\right)$ (14)
+$+$ $\sum_{j\in\Lambda_{2}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+b_{j}\right)-f(\mathbf{x})|<\varepsilon$
+Since $|\sum_{j\in\Lambda_{2}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)|\to 0$, we have the following inequality:
 $\big{|}|\sum_{j\in\Lambda_{1}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)-f(\mathbf{x})\mid$ (15)
@@
 Therefore, we have:
-$|\sum_{j=1}^{N_{1}}\alpha_{j}\sigma\left(\mathbf{W}j^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)-f(\mathbf{x})|$ (15)
-$<\varepsilon+|\sum_{j=1}^{N_{2}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)|$
+$|\sum_{j\in\Lambda_{1}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+b_{j}\right)-f(\mathbf{x})\mid$ (16)
+$<$ $\varepsilon+|\sum_{j\in\Lambda_{2}}\alpha_{j}\sigma\left(\mathbf{W}j^{\mathrm{T}}\mathbf{x}+b_{j}\right)|$
 Hence, when parameters in certain layers are small enough, we can directly remove those layers since their impact on the final result is minimal.
@@
 Re-training models from scratch incurs significant computational costs, so reusing well-trained models for new tasks is both practical and resource-efficient.
-A prominent solution to this challenge is the LoRA (Hu et al.
-2021), which can be expressed as follows:
+A prominent solution to this challenge is the LoRA Hu et al.
+(2021), which can be expressed as follows:
 $\mathbf{h}=\mathbf{W}_{0}\mathbf{x}+\Delta\mathbf{W}\mathbf{x}$ (17)
@@
 $\mathbf{x}_{i+1}$ $=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,1})+\sum_{j=1}^{i+1}(\mathbf{W}^{\prime}_{j,3}+\Delta\mathbf{W}_{j+1,3})$ (18)
-$\sigma((\mathbf{W}^{\prime}_{j,2}+\Delta\mathbf{W}_{j+1,2})\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$
-From Eq. 17, it can be seen that LoRA essentially fine-tunes the UAT parameters layer by layer for a specific task.
-Figure 7: The comparison between Human and LLMs.
-4.5 Rethinking LLMs
+$+\sigma((\mathbf{W}^{\prime}_{j,2}+\Delta\mathbf{W}_{j+1,2})\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$
+From Eq. 18, it can be seen that LoRA essentially fine-tunes the DUAT parameters layer by layer for a specific task.
+5 Conclusion
+In this paper, we delve into the theoretical underpinnings of LLMs, demonstrating that contemporary LLMs, primarily constructed with Transformer architectures, embody concrete manifestations of the DUAT.
+The remarkable generalization prowess exhibited by LLMs is attributed to their MHA modules and residual operation, which enable the adaptation to approximate diverse functions based on the presented input data.
+Contextual interaction emerges as a paramount capability for LLMs, manifesting such abilities as ICL, instruction following, and contextual reasoning.
+These competencies are enabled by the Transformer’s MHA to learn from context.
+Expanding upon this understanding, we have provided a rigorous theoretical grounding for key techniques employed in LLMs, including LoRA for efficient fine-tuning and pruning for model compression, elucidating their effectiveness through the lens of the UAT.
+By leveraging the theoretical framework provided by the DUAT, not only can existing methodologies be explained but also avenues for the future evolution of LLMs are illuminated.
+6 Rethinking LLMs
 The capabilities of LLMs have become so advanced that their language processing abilities are approaching human levels, raising a core question: How do LLMs differ from humans in language processing?
@@
 At this level, there is almost no difference.
-Given the ambiguity of words, determining context is crucial: humans understand context through the activation and transmission of neurons in the brain, while LLMs approximate the corresponding functions using UAT.
+Given the ambiguity of words, determining context is crucial: humans understand context through the activation and transmission of neurons in the brain, while LLMs approximate the corresponding functions using DUAT.
 Here, the input and output of network layers are analogous to the transmission of neural signals in the brain, and the function approximation corresponds to the final output of humans.
@@
 Are these also capabilities of LLMs?
-First, we can consider the human brain as a combination of one or more UAT models, which are randomly initialized at birth.
+First, we can consider the human brain as a combination of one or more DUAT models, which are randomly initialized at birth.
 What supports such an assumption?
-The original neural network, the perceptron (Rosenblatt 1963), was designed based on the human neuron.
+The original neural network, the perceptron Rosenblatt (1963), was designed based on the human neuron.
 Over time, as deep learning networks evolved, the difference in form between deep networks and the perceptron grew, leading people to no longer associate neural networks with human neurons.
-However, through our derivation, the Transformer can also be understood as an implementation of UAT, and the perceptron can be seen as a layer of UAT.
-The human brain can be considered as a cluster of multi-layer perception.
-Learning language in a social environment is akin to training the UATs in the brain.
+However, through our derivation, the Transformer can also be understood as an implementation of DUAT, and the perceptron can be seen as a simple UAT.
+So the human brain can be considered as a cluster of multi-layer perception.
+Learning language in a social environment is akin to training the DUATs in the brain.
 Therefore, what we call understanding, analysis, and memory retrieval are essentially processes of fitting outputs based on inputs.
-Understanding and analysis: When we respond to a question posed by someone, it can be seen as a simple example of understanding and analysis, which both humans and LLMs can accomplish.
-Humans may guess the answer to unfamiliar questions based on their experiences, which can be either correct or incorrect, and LLMs exhibit similar behavior.
-Memory retrieval: This is essentially fitting specific results based on particular words, as there is no actual database in the brain.
+Taking memory retrieval as an example: This is essentially fitting specific results based on particular words, as there is no actual database in the brain.
 For instance, recalling the experience of eating an apple for the first time in childhood is the brain fitting specific results based on those words; without those words, the memory would not surface.
 This is because the brain dynamically outputs results based on inputs and fills in the details of the event based on learned natural rules (We provide some examples about understanding, analysis of language, and memory retrieval in Appendix E).
-However, as humans grow, the weights in the brain are constantly updated, leading to potential memory distortions.
+However, as humans grow, the weights in the brain are constantly updated, leading to potential memory distortions (Wang & Li, 2024a).
 Therefore, it is entirely reasonable for LLMs to make errors or generate hallucinations—these are just outputs produced based on existing weights and inputs, a problem humans also face.
-We believe the greatest advantage humans have over LLMs is their powerful multimodal and multitask processing abilities (which can be understood as the coordination of multiple UAT models to produce reasonable results).
+We believe the greatest advantage humans have over LLMs is their powerful multimodal and multitask processing abilities (which can be understood as the coordination of multiple DUAT models to produce reasonable results).
 Another advantage is the interaction with the real world, which allows us to verify the knowledge we’ve learned in reality, enabling the brain to continuously optimize its parameters based on inputs.
 LLMs, on the other hand, are limited to function approximation within the corpora data.
-5 Conclusion
-In this paper, we delve into the theoretical underpinnings of LLMs, demonstrating that contemporary LLMs, primarily constructed with Transformer architectures, embody concrete manifestations of the UAT.
-The remarkable generalization prowess exhibited by LLMs is attributed to their MHA modules and residual operation, which enable the adaptation to approximate diverse functions based on the presented input data.
-Contextual interaction emerges as a paramount capability for LLMs, manifesting such abilities as ICL, instruction following, and contextual reasoning.
-These competencies are enabled by the Transformer’s MHA to learn from context.
-Expanding upon this understanding, we have provided a rigorous theoretical grounding for key techniques employed in LLMs, including LoRA for efficient fine-tuning and pruning for model compression, elucidating their effectiveness through the lens of the UAT.
-By leveraging the theoretical framework provided by the UAT, not only can existing methodologies be explained but also avenues for the future evolution of LLMs are illuminated.
-Appendix A The Property of Diamond
-[…]
-$\mathbf{W}_{1}\diamond[\mathbf{x}\diamond\mathbf{W}_{2}]$ (1)
-$=$ $\mathbf{W}_{1}\diamond[\mathbf{W}_{2}^{T}\mathbf{x}]$
-$=$ $\mathbf{W}_{1}^{T}[\mathbf{W}_{2}^{T}\mathbf{x}]$
-$=$ $\mathbf{W}_{1}^{T}\mathbf{W}_{2}^{T}\mathbf{x}=\mathbf{W}_{12}\mathbf{x}$
-$=$ $(\mathbf{W}_{1}^{T}\mathbf{W}_{2}^{T})^{T}\diamond\mathbf{x}$
-$=$ $(\mathbf{W}_{2}\mathbf{W}_{1})\diamond\mathbf{x}$
-$=$ $\mathbf{W}_{2}^{T}\diamond\mathbf{W}_{1}\diamond\mathbf{x}$
-$[\mathbf{W}_{1}\diamond\mathbf{x}]\diamond\mathbf{W}_{2}$ (2)
-$=$ $[\mathbf{W}_{1}^{T}\mathbf{x}]\diamond\mathbf{W}_{2}$
-$=$ $\mathbf{W}_{2}^{T}[\mathbf{W}_{1}^{T}\mathbf{x}]$
-$=$ $[\mathbf{W}_{2}^{T}\mathbf{W}_{1}^{T}]^{T}\diamond\mathbf{x}$
-$=$ $[\mathbf{W}_{1}\mathbf{W}_{2}]\diamond\mathbf{x}$
-$=$ $\mathbf{W}_{1}^{T}\diamond\mathbf{W}_{2}\diamond\mathbf{x}$
-Figure 1: The process of transforming $\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}}_{8})$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{HV}^{\prime}\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$.
-Figure 2: The process of Transformer.
-Appendix B The Matrix-Vector Form of MHA
+Figure 7: The comparison between Human and LLMs.
+Appendix A The Matrix-Vector Form of MHA
 In this section, we will present the transformation process of MHA into their corresponding matrix-vector forms.
 Figure 1.a represents the computation process of obtaining $\mathbf{W}_{QKi}$ in the attention mechanism, while Figure 1.b computes $\mathbf{H}_{1},\dots,\mathbf{H}_{8}$ through the softmax operation.
-Figure 1.c and Figure 1.e illustrates the process of converting $\operatorname{Concat}\left(\mathbf{H}_{1}\mathbf{V}_{1},\ldots,\mathbf{H}_{h}\mathbf{V}_{h}\right)$ into the matrix-vector form $\mathbf{W}_{HV}^{\prime}\diamond\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{HV}^{\prime}$ is generated from $\mathbf{H}_{1},\dots,\mathbf{H}_{8}$ and $\mathbf{W}_{V1},\dots,\mathbf{W}_{V8}$.
+Figure 1.c and Figure 1.e illustrates the process of converting $\operatorname{Concat}\left(\mathbf{H}_{1}\mathbf{V}_{1},\ldots,\mathbf{H}_{h}\mathbf{V}_{h}\right)$ into the matrix-vector form $(\mathbf{W}_{HV}^{\prime})^{T}\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{HV}^{\prime}$ is generated from $\mathbf{H}_{1},\dots,\mathbf{H}_{8}$ and $\mathbf{W}_{V1},\dots,\mathbf{W}_{V8}$.
 Figure 1.c shows $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$.
-In Figure 1.d, we provide a simple example demonstrating the conversion of $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$ into $\mathbf{W}_{HVi}^{\prime}\mathbf{x}_{i}^{\prime}$.
+In Figure 1.d, we provide a simple example demonstrating the conversion of $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$ into $(\mathbf{W}_{HVi}^{\prime})^{T}\mathbf{x}_{i}^{\prime}$.
 This process is divided into four parts: Figure 1.d.1 represents the general form of $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$, while Figure 1.d.2 serves as a simple example of Figure 1.d.1.
-Figure 1.d.3 first rewrites Figure 1.d.2 using diamond multiplication as $\mathbf{H}_{i}^{\prime}\diamond[\mathbf{x}_{i}^{\prime}\diamond\mathbf{W}_{Vi}^{\prime}]$, and then utilizes the property of diamond multiplication to express it as $(\mathbf{W}_{Vi}^{\prime})^{T}\diamond\mathbf{H}_{i}^{\prime}\diamond\mathbf{x}_{i}^{\prime}$ in Figure 1.d.4.
-Finally, Figure 1.d.5 is obtained as $\mathbf{W}_{HVi}^{\prime}\diamond\mathbf{x}_{i}^{\prime}=\mathbf{H}_{i}^{\prime}$.
+Figure 1.d.3 first rewrites Figure 1.d.2 using matrix multiplication as $[\mathbf{H}_{i}^{\prime}\mathbf{W}_{Vi}^{\prime})]^{T}\mathbf{x}_{i}^{\prime}=\hat{\mathbf{H}}_{i}^{\prime}$, and then express it as $(\mathbf{H}_{i}^{\prime})^{T}[\mathbf{W}_{Vi}^{\prime})^{T}\mathbf{x}_{i}^{\prime}]=\hat{\mathbf{H}}_{i}^{\prime}$ in Figure 1.d.4.
+Finally, Figure 1.d.5 is obtained as $(\mathbf{W}_{HVi}^{\prime})^{T}\mathbf{x}_{i}^{\prime}=(\hat{\mathbf{H}}_{i}^{\prime})^{T}$.
 It can be observed that $\mathbf{W}_{HVi}^{\prime}$ is a dense matrix.
-In Figure 1.e, we present a simple example of $\mathbf{H}_{1}[\mathbf{x}_{1}\mathbf{W}_{V1}],\mathbf{H}_{2}[\mathbf{x}_{2}\mathbf{W}_{V2}]...$ into the matrix-vector form $\mathbf{W}_{HV}^{\prime}\mathbf{x}^{\prime}$.
+In Figure 1.e, we present a simple example of $\mathbf{H}_{1}[\mathbf{x}_{1}\mathbf{W}_{V1}],\mathbf{H}_{2}[\mathbf{x}_{2}\mathbf{W}_{V2}]...$ into the matrix-vector form $(\mathbf{W}_{HV}^{\prime})^{T}\mathbf{x}^{\prime}=(\hat{\mathbf{H}})^{T}$.
 Figure 1.e.1 depicts $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$, while Figure 1.e.2 represents the Matrix-vector format of Figure 1.e.1.
-Appendix C The UAT Format of Transformer
-In this section, we will prove multi-layer Transformer is the implementation of UAT.
-C.1 The Properties of UAT
-Before expressing Transformers in the UAT format, we present two lemma regarding UAT.
+Figure 1: The process of transforming $\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}}_{8})$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{HV}^{\prime}\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$.
+Appendix B Transformer: The implementation of DUAT
+In this section, we will prove multi-layer Transformer is the implementation of DUAT.
+B.1 The Properties of UAT
+Before expressing residual-based CNNs and Transformer-based ViTs in the UAT format, we present a lemma regarding UAT.
 There are two cases for UAT-approximated functions: $f(\mathbf{x})\in\mathbb{R}$ and $f(\mathbf{x})\in\mathbb{R}^{m}$.
@@
 Lemma 1. The mathematical form of UAT remains unchanged when multiplied by a matrix (constant).
-$G(\mathbf{x})$ $=\mathbf{\beta}\sum_{j=1}^{N}\mathbf{\alpha}_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\mathbf{\theta}_{j}\right)$ (3)
-$=\sum_{j=1}^{N}\mathbf{\beta}\mathbf{\alpha}_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\mathbf{\theta}_{j}\right)$
+$G(\mathbf{x})$ $=\mathbf{\beta}\sum_{j=1}^{N}\mathbf{\alpha}_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\mathbf{b}_{j}\right)$ (19)
+$=\sum_{j=1}^{N}\mathbf{\beta}\mathbf{\alpha}_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\mathbf{b}_{j}\right)$
 Eq.
-(3) shows the representation of UAT multiplying a matrix.
-Let $\mathbf{\alpha}_{j}=\mathbf{\beta}\mathbf{\alpha}_{j}$, and the result remains consistent with the original UAT mathematical form.
+19 shows the representation of UAT multiplying a matrix.
+Let $\mathbf{\alpha}_{j}=\mathbf{\beta}\mathbf{\alpha}_{j}$, and the general mathematical form of Eq.
+19 remains consistent with the original UAT mathematical form.
 Thus, it is proven that the mathematical form of UAT remains unchanged when multiplied by a matrix (constant).
-Lemma 2.
-Let the general term of the network be written as: $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, $i=0,1,2,3,\ldots$.
-The multi-layer network of this form still corresponds to the mathematical form of UAT, and we refer to such a network as a residual network.
-To prove that a multi-layer network with the general term written as $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$ corresponds to the UAT mathematical form, we first provide the forms of single-layer and two-layer networks, as shown in Eq.
-(4) and (5).
+B.2 From UAT to DUAT
+In this section, we first present a general form for a single-layer residual term in a network and demonstrate that the mathematical form of a multi-layer network composed of this residual term aligns with the DUAT framework.
+Before proceeding with the formal proof, we first define DUAT explicitly: DUAT shares the overall mathematical form of UAT, but with certain parameters influenced by the input, allowing them to change dynamically in response to it.
+We refer to these parameters as ”dynamic parameters.” In standard UAT, parameters remain fixed once training is complete, whereas, in DUAT, the dynamic parameters are functions of the input and thus vary with it.
+Consequently, dynamic parameters in DUAT may take the form of complex functions.
+Currently, DUAT primarily is implemented by residual structure, meaning that, in general, these complex functions are also DUAT.
+Next, we will give the proof.
+A general residual term of the network can be written as:
+$\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i-1}+\mathbf{b}^{\prime}_{i,3})+\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i,2}\mathbf{x}^{\prime}_{i-1}+\mathbf{b}^{\prime}_{i,2})$ (20)
+where $i=1,2,3,\ldots$ and $\mathbf{x}^{\prime}_{0}$ represents the input.
+A multi-layer network structured in this way aligns with the mathematical form of DUAT.
+To demonstrate that a multi-layer network leverages the general term of Eq.
+20 corresponds to DUAT, we first examine the forms of single-layer and two-layer networks, as shown in Eqs.
+21 and 22.
 $\mathbf{x}^{\prime}_{1}=(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ (21)
@@
 In Eq.
-(5), let $\mathbf{W}^{\prime}_{2,1}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,1}}$, $\mathbf{b}^{\prime}_{2,3}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,3}}$, $\mathbf{W}^{\prime}_{1,3}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}}$, $\mathbf{W}^{\prime}_{2,3}=\underline{\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}}$, and $\mathbf{b}^{\prime}_{2,2}=\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}$.
+22, let $\mathbf{W}^{\prime}_{2,1}=\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,1}$, $\mathbf{b}^{\prime}_{2,1}=\mathbf{W}^{\prime}_{2,1}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,3}$, $\mathbf{W}^{\prime}_{1,3}=\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}$, $\mathbf{W}^{\prime}_{2,2}=\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}$, and $\mathbf{b}^{\prime}_{2,2}=(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$.
 Thus, Eq.
-(6) can be written as:
+22 can be written into Eq.
+23.
 $\mathbf{x}^{\prime}_{2}$ $=(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,1})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ (23)
 $+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,2})$
-Why can we combine the above terms?
-Once the network training is complete, we know these parameters, so they can be directly calculated and thus combined.
-And $\mathbf{b}^{\prime}_{2,2}=\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}$ clearly has the same mathematical form as UAT.
-We can understand it as dynamically fitting the bias term $\mathbf{b}^{\prime}_{2,2}$ using UAT based on the input.
-Therefore, we have proven that the mathematical form of one and two-layer residual networks is consistent with the UAT mathematical form.
-Assume that the mathematical form of the first $i$ layers of the residual network is consistent with UAT.
-Our goal is to prove that the mathematical form of the $i+1$-th layer of the residual network is still consistent with UAT.
-For the convenience of expression, we make the following definition: since the mathematical form of the first $i$ layers of the residual network is consistent with UAT, we write $\mathbf{x}^{\prime}_{i}$ as $\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,3})+UAT^{R}_{i}$, where the first term is written separately and the rest is written as the remainder term $UAT^{R}_{i}$.
+For the cleatity, we could write Eq.
+21 and Eq.
+23 into Eq.
+24 and Eq.
+25, where $UAT^{R}_{1}=\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ and $UAT^{R}_{2}=\Sigma_{j=1}^{2}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$.
+$\mathbf{x}^{\prime}_{1}=(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+UAT^{R}_{1}$ (24)
+$\mathbf{x}^{\prime}_{2}$ $=(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,1})+UAT^{R}_{2}$ (25)
+According to Eq.
+24 and Eq.
+25, aside from the initial term, the overall mathematical forms of the residual terms $UAT^{R}_{1}$ and $UAT^{R}_{2}$ are consistent with UAT.
+However, for $UAT^{R}_{2}$, its parameter $\mathbf{b}^{\prime}_{2,2}$ is influenced by the input, while other weight parameters, such as $\mathbf{W}^{\prime}_{2,1}$, can be disregarded.
+These parameters are only affected by other parameters, so once the network training is complete, they are essentially fixed.
+This allows us to focus primarily on the dynamic parameters.
+In equation $\mathbf{b}^{\prime}_{2,2}=(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$, except for the initial term, $\mathbf{b}^{\prime}_{2,2}$ also follows the same mathematical form as UAT.
+This can be interpreted as dynamically adjusting the bias term $\mathbf{b}^{\prime}_{2,2}$ via UAT based on the input.
+In conclusion, we have demonstrated that the mathematical forms of one-layer and two-layer residual networks are consistent with the UAT framework and the two-layer residual network is the DUAT function.
+Next, we will use mathematical induction to prove that the mathematical form of a multi-layer residual network is also a DUAT function.
+Assume that the overall mathematical form of the first $i$ layers of the residual network aligns with UAT.
+Our goal is to show that the overall mathematical form of the $i+1$-th layer remains consistent with UAT.
+For clarity, we make the following definitions: since the overall mathematical form of the first $i$ layers is consistent with UAT, we can write it as $\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,1})+UAT^{R}_{i}$, where the first term is explicitly written, and the remainder is denoted as $UAT^{R}_{i}$, with $UAT_{i}^{R}=\sum_{j=1}^{i}\mathbf{W}^{\prime}_{j,2}\sigma(\mathbf{W}^{\prime}_{j,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,1})$.
 Since $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, we divide $\mathbf{x}^{\prime}_{i+1}$ into two parts: $(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})$ and $\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$.
-First, consider the part $(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})$.
-Substitute $\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,3})+UAT^{R}_{i}$ into it, and we get Eq.
-(7).
-Let $\mathbf{W}^{\prime}_{i+1,1}=\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}^{\prime}_{i,1}}$ and $\mathbf{b}^{\prime}_{i+1,3}=\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{b}^{\prime}_{i,1}+\mathbf{b}^{\prime}_{i+1,3}}$, we can simplify the first part to $(\mathbf{W}^{\prime}_{i+1,}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,1}UAT^{R}_{i}$.
-Since the mathematical form of $UAT^{R}_{i}$ is consistent with UAT, and we have already proven that the mathematical form of UAT remains unchanged when multiplied by a matrix, we have proven that the mathematical form of the first part is consistent with UAT.
+First, consider $(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})$.
+Substituting $\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,1})+UAT^{R}_{i}$, we obtain Eq.
+26.
+Setting $\mathbf{W}^{\prime}_{i+1,1}=\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}^{\prime}_{i,1}$ and $\mathbf{b}^{\prime}_{i+1,1}=\mathbf{W}^{\prime}_{i+1,1}\mathbf{b}^{\prime}_{i,1}+\mathbf{b}^{\prime}_{i+1,3}$, we can simplify the first part to $(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i+1,1})+\mathbf{W}^{\prime}_{i+1,1}UAT^{R}_{i}$.
+Since the overall mathematical form of $UAT^{R}_{i}$ is consistent with the $i$-layer UAT and we have shown that the UAT form is preserved when multiplied by a matrix, we have thus demonstrated that the overall mathematical form of the first part remains consistent with UAT.
 $(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})$ (26)
@@
 $=$ $\{[\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{0}+\underline{(\mathbf{W}^{\prime}_{i+1,1}\mathbf{b}^{\prime}_{i,1}+\mathbf{b}^{\prime}_{i+1,3})}]+\mathbf{W}^{\prime}_{i+1,1}UAT^{R}_{i}\}$
-Next, we prove that the mathematical form of the second part, $\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, is consistent with UAT.
-By substituting $\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,3})+UAT^{R}_{i}$ into it, we obtain Eq.
-(8).
-Let $\mathbf{W}^{\prime}_{i+1,2}=\underline{\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}^{\prime}_{i,1}}$ and $\mathbf{b}^{\prime}_{i+1,2}=\underline{(\mathbf{W}^{\prime}_{i+1,2}\mathbf{b}^{\prime}_{i,1}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{W}^{\prime}_{i+1,2}UAT^{R}_{i}}$.
-Thus, the second term can be written as $\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i+1,2})$, which can be considered as a term within UAT.
-Here, $\mathbf{b}^{\prime}_{i+1,2}$ is a bias term approximated using UAT.
+Next, we show that the overall mathematical form of the second part, $\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, is also consistent with UAT.
+Substituting $\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,1})+UAT^{R}_{i}$ into this expression, we arrive at Eq.
+27.
+Setting $\mathbf{W}^{\prime}_{i+1,2}=\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}^{\prime}_{i,1}$ and $\mathbf{b}^{\prime}_{i+1,2}=(\mathbf{W}^{\prime}_{i+1,2}\mathbf{b}^{\prime}_{i,1}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{W}^{\prime}_{i+1,2}UAT^{R}_{i}$, we can rewrite the second part as $\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i+1,2})$, which can be interpreted as a term in UAT.
+Here, $\mathbf{b}^{\prime}_{i+1,2}$ serves as a bias term approximated using DUAT.
 $\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$ (27)
@@
 $=$ $\mathbf{W}^{\prime}_{i+1,3}\sigma\{\underline{\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{0}+\underline{(\mathbf{W}^{\prime}_{i+1,2}\mathbf{b}^{\prime}_{i,1}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{W}^{\prime}_{i+1,2}UAT^{R}_{i}}\}$
-Since the first term of $\mathbf{x}^{\prime}_{i+1}$ has the same mathematical form as UAT, and the second term can be considered as a component within UAT, the sum of both terms still corresponds to the UAT mathematical form.
-Therefore, we have proven that the mathematical form of a multi-layer residual network is consistent with the UAT mathematical form.
-The difference lies in that some bias terms are approximated using UAT rather than being directly defined.
-C.2 Transformer to UAT
-Similarly, To illustrate relationship between multi-layer Transformer and UAT, we will first derive the general form based on Figure 2 and then compare it with the general form given in Section C.1: $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2}).$
-By examining the relationship between these forms, we can demonstrate that the mathematical form of multi-layer Transformer networks is consistent with the UAT.
-The Transformer involves two key operations: MHA and FFN.
-These operations, when expressed in matrix-vector form, correspond to Eq.(10) and Eq.(6), respectively.
-Consequently, the general term for a Transformer-based network can be written as Eq.
-(9).
+Let $UAT^{R}_{i+1}=\mathbf{W}^{\prime}_{i+1,1}UAT^{R}_{i}+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i+1,2})$.
+Then, we can express $\mathbf{x}^{\prime}_{i+1}$ as $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i+1,1})+UAT^{R}_{i+1}$.
+Thus, apart from the initial term, the overall mathematical form of $\mathbf{x}^{\prime}_{i+1}$ remains consistent with UAT.
+Since a single term does not alter the nature of UAT, we conclude that the overall form of a multi-layer residual network aligns with UAT.
+Furthermore, as certain bias parameters within the network, such as $\mathbf{b}^{\prime}_{i+1,2}$, are also approximated using UAT, we have therefore demonstrated that the mathematical form of a multi-layer residual network is a UAT function.
+B.3 Transformer to DUAT
+Similarly, to prove that Transformer-based ViTs are also DUAT functions, we start by deriving their general form based on Figure 9 in the main text.
+Transformers involve two key operations: MHA and the Feed-Forward Network (FFN).
+In matrix-vector form, these operations correspond to Equations 28 and 29.
+$MHA(\mathbf{x}_{i})\mapsto\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}$ (28)
+$FFN(\mathbf{x}_{i})\mapsto\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}\$ (29)
 $\mathbf{x}^{\prime}_{i+1}$ $=\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i+1,3}\sigma[\mathbf{W}^{\prime}_{i+1,2}(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i})+\mathbf{b}^{\prime}_{i+1,2}]+\mathbf{b}^{\prime}_{i+1,3}$ (30)
 $=\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i+1,3}\sigma(\underline{\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}^{\prime}_{i+1,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{b}^{\prime}_{i+1,3}$
-Let $\mathbf{W}^{\prime}_{i+1,2}=\underline{\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}^{\prime}_{i+1,1}}$, so we have $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$.
-It is evident that, when compared to the general form given in Section C.1: $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, the mathematical forms are the same.
-Therefore, multi-layer Transformers are also specific implementations of the UAT.
-Similarly, to clearly present the UAT format, we use the method mentioned in Section C.1.
-Assuming the mathematical form of $\mathbf{x}^{\prime}_{i}$ is consistent with UAT, we decompose $\mathbf{x}^{\prime}_{i}$ into a primary term plus a remainder term, written as $\mathbf{x}^{\prime}_{i}=(\mathbf{W}_{i,1}\mathbf{x}_{0}+\mathbf{b}_{i,3})+UAT^{R}_{i}$.
-Substituting this into $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, we obtain Eq.
-(11).
-By setting $\mathbf{W}^{\prime}_{i+1,1}=\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}_{i,1}}$, $\mathbf{b}^{\prime}_{i+1,3}=\underline{(\mathbf{W}^{\prime}_{i+1,1}\mathbf{b}_{i,1}+\mathbf{b}^{\prime}_{i+1,3})}$, $UAT^{R}_{i}=\mathbf{W}^{\prime}_{i+1,1}UAT^{R}_{i}$, and $\mathbf{W}^{\prime}_{i+1,2}=\underline{\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}_{i,1}}$, $\mathbf{b}^{\prime}_{i+1,2}=\underline{(\mathbf{W}^{\prime}_{i+1,2}\mathbf{b}_{i,1}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{W}^{\prime}_{i+1,2}UAT^{R}_{i}}$, we get $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}_{0}+\mathbf{b}^{\prime}_{i+1,3})+UAT^{R}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}_{0}+\mathbf{b}^{\prime}_{i+1,2})$.
-Since the mathematical form of $UAT^{R}_{i}$ is consistent with multi-layer UAT, it follows that multi-layer Transformer networks are also specific implementations of UAT.
-In addition, $UAT^{R}_{i+1}=UAT^{R}_{i}+\mathbf{W}_{i+1,3}^{\prime}\sigma(\mathbf{W}_{i+1,2}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,2})$.
-So, $\mathbf{x}_{i+1}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,3})+UAT^{R}_{i+1}$, where $i=0,1,2\ldots$, $UAT_{i+1}^{R}=\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$, $\mathbf{b}^{\prime}_{j,2}=(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})+\mathbf{W}^{\prime}_{j,2}UAT_{j-1}^{R}$.
-To sum up, the $i+1$-layer Transformer can be written as:
-$\mathbf{x}_{i+1}^{\prime}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}^{\prime}+\mathbf{b}_{i+1,1})+\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$ (10)
-Other parameters will also vary with the number of layers.
-And, these changes involve only matrix multiplications or additions.
-Thus, they do not fundamentally alter the mathematical form of the UAT.
-Therefore, we will not detail the specific forms for each layer here.
-The parameters in the MHA mechanism dynamically change with the input.
-Therefore, in the corresponding UAT mathematical form for Transformers, all $\mathbf{W}^{\prime}_{i,1}$, $\mathbf{W}^{\prime}_{j,1}$ and $\mathbf{W}^{\prime}_{j,2}$ parameters for the $i$-th layer, where $j=1,...,i$, are dynamically changing with the input.
-We give some examples in Figure 3.
+Thus, a general term in a Transformer-based network can be expressed as Equation 30.
+Letting $\mathbf{W}^{\prime}_{i+1,2}=\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}^{\prime}_{i+1,1}$, we can rewrite the general term as follows:
+$\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$ (31)
+Clearly, Eq.
+31 matches the mathematical form presented in Section B.2: $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$.
+Therefore, a multilayer Transformer-based ViT is indeed DUAT function.
+To clearly show the DUAT format of a multilayer ViT, we use the approach outlined in Section B.2.
+Assuming that the overall mathematical form of $\mathbf{x}^{\prime}_{i}$ aligns with the UAT structure, we decompose $\mathbf{x}^{\prime}_{i}$ into a main term and a residual term, expressed as $\mathbf{x}^{\prime}_{i}=(\mathbf{W}_{i,1}\mathbf{x}_{0}+\mathbf{b}_{i,3})+UAT^{R}_{i}$, where $UAT_{i}^{R}=\sum_{j=1}^{i}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$ and $i=1,2,\ldots$.
+Substituting this into $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, we get Eq.
+32.
 $\mathbf{x}^{\prime}_{i+1}$ $=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$ (32)
@@
 $+\mathbf{W}^{\prime}_{i,3}\sigma\{\underline{\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}_{i,1}}\mathbf{x}_{0}+\underline{(\mathbf{W}^{\prime}_{i+1,2}\mathbf{b}_{i,1}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{W}^{\prime}_{i+1,2}UAT^{R}_{i}}\}$
-Figure 3: Some examples of the UAT format of multi-layer Transformer.
+Let $\mathbf{W}^{\prime}_{i+1,1}=\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}_{i,1}$, $\mathbf{b}^{\prime}_{i+1,3}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{b}_{i,1}+\mathbf{b}^{\prime}_{i+1,3})$, $UAT^{R}_{i}=\mathbf{W}^{\prime}_{i+1,1}UAT^{R}_{i}$, $\mathbf{W}^{\prime}_{i+1,2}=\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}_{i,1}$, and $\mathbf{b}^{\prime}_{i+1,2}=(\mathbf{W}^{\prime}_{i+1,2}\mathbf{b}_{i,1}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{W}^{\prime}_{i+1,2}UAT^{R}_{i}$.
+Substituting these, we obtain $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}_{0}+\mathbf{b}^{\prime}_{i+1,3})+UAT^{R}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}_{0}+\mathbf{b}^{\prime}_{i+1,2})$.
+Since $UAT^{R}_{i}$ overall aligns with the UAT, the overall mathematical form of the $i+1$ layers’ ViT also conforms to the UAT.
+So the $i+1$ layers’ ViT can be expressed as:
+$\mathbf{x}_{i+1}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,3})+\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$ (33)
+When $i=0$, the parameters are defined as follows: $\mathbf{W}^{\prime}_{1,1}=\mathbf{W}^{\prime}_{1,1}$, $\mathbf{b}^{\prime}_{1,3}=\mathbf{b}^{\prime}_{1,3}$, $\mathbf{W}^{\prime}_{1,3}=\mathbf{W}^{\prime}_{1,3}$, $\mathbf{W}^{\prime}_{1,2}=\mathbf{W}^{\prime}_{1,2}\mathbf{W}^{\prime}_{1,1}$, and $\mathbf{b}^{\prime}_{1,2}=\mathbf{b}^{\prime}_{1,2}$.
+We define the parameters for cases where $i\geq 1$.
+For each $j=1,2,\ldots,i$, the updates are as follows: $\mathbf{W}^{\prime}_{j+1,1}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{W}_{j,1}$, $\mathbf{b}^{\prime}_{j+1,3}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{b}_{j,3}+\mathbf{b}^{\prime}_{j+1,3}$, $\mathbf{W}^{\prime}_{j+1,2}=\mathbf{W}^{\prime}_{j+1,2}\mathbf{W}_{j,1}$, and $\mathbf{W}^{\prime}_{j,3}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{W}^{\prime}_{j,3}$.
+Additionally, for $j=1,2,\ldots,i+1$, the bias terms $\mathbf{b}^{\prime}_{j,2}$ are updated according to:
+$\mathbf{b}^{\prime}_{j,2}=$ $(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})$ (34)
+$+$ $\mathbf{W}^{\prime}_{j,2}\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,3}\sigma(\mathbf{W}^{\prime}_{k,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,2}).$
+It means that the whole computing process is serial, if we want to compute $\mathbf{W}^{\prime}_{3,1}$, we must calculate $\mathbf{W}^{\prime}_{2,1}$, because $\mathbf{W}^{\prime}_{3,1}=\mathbf{W}^{\prime}_{3,1}\mathbf{W}^{\prime}_{2,1}$.
+Thus, we have established that a multilayer ViT is also a DUAT function.
+The mathematical representation of the ViT aligns with the DUAT framework, with $\mathbf{b}^{\prime}_{j,2}$ closely approximated by DUAT.
+The most notable difference between ViTs and residual-based CNNs is the dynamic input-dependent nature of the parameters in the MHA mechanism.
+Consequently, in the DUAT formulation associated with ViT, the parameters $\mathbf{W}^{\prime}_{j,1}$ and $\mathbf{W}^{\prime}_{j,2}$ for $i+1\geq j\geq 1$, and $\mathbf{W}^{\prime}_{j,3}$ for $i+1>j>1$ in layer $i$, all adapt dynamically based on the input.Besides, we provide some examples in Figure 2.
+Figure 2: Some examples of the DUAT format of multi-layer Transformer.
 The changes of parameters are represented within the dashed boxes.
@@
 There is no specific order of calculation for the parameters within each dashed box, but there is a top-to-bottom calculation order between different dashed boxes.
-Appendix D The Learning Process of MHA
+Appendix C The Learning Process of MHA
+In this section, we will present the specific mathematical form of the parameters $\mathbf{H}_{i}$ in the MHA shown in Figure 1.
+First, it can be written as Eq.
+35.
+Then, in Figure 3, we illustrate how each parameter in $\mathbf{H}_{i}$ is calculated.
+It is clear that these parameters are learned elements of the covariance matrix of the input.
+This process disrupts the overall mathematical form of the UAT, so $\mathbf{H}_{i}$ can be simply and directly regarded as parameters learned from the covariance matrix of the upper-layer input.
+$\mathbf{H}_{i}=$ $\operatorname{softmax}(\frac{\mathbf{x}_{i}\mathbf{W}_{Qi}(\mathbf{x}_{i}\mathbf{W}_{Ki})^{T})}{\sqrt{M}})$ (35)
+$=$ $\operatorname{softmax}(\frac{\mathbf{x}_{i}\mathbf{W}_{Qi}\mathbf{W}_{Ki}^{T}\mathbf{x}_{i}^{T}}{\sqrt{M}})$
+$=$ $\operatorname{softmax}(\frac{\mathbf{x}_{i}\mathbf{W}_{QKi}\mathbf{x}_{i}^{T}}{\sqrt{M}})$
+Figure 3: The transformation process of the $\mathbf{H}_{i}$.
 To more clearly illustrate the learning details of the MHA mechanism in Transformers, we present an example in Figure 4.
@@
 Figure 5: The learning process of 1D convolution.
-Appendix E The Comparision between Human and LLMs
+Appendix D The Comparision between Human and LLMs
 To more specifically compare the differences between humans and LLMs, we provide a simple example here.
@@
 relativity theory → Albert Einstein.” This process can be understood as understanding the question, analyzing it, retrieving memories, and producing results.
-However, if the human brain is considered as a UAT, it generates the output based on specific inputs, such as ”gravitation → Isaac Newton” or relevant information about the theory itself.
+However, if the human brain is considered as a DUAT, it generates the output based on specific inputs, such as ”gravitation → Isaac Newton” or relevant information about the theory itself.
 But the truth is that the brain never creates a ’memory warehouse’ to store that information, it just produces the output based on input.

### v5

@@
 where $\mathbf{W}=\mathbf{W}_{1}\cdots\mathbf{W}_{N}$ and $\mathbf{b}=b_{1}\cdots b_{N}$igure 2 illustrates the comparison between UAT and DUAT.
-On the left side, Figure 2.a demonstrates the deficiency in data recognition.
-The input to the network is shown, with only the initial part differing slightly.
-This can be likened to tasks in LLMs, such as translating, continuing, or summarizing the same passage, where only the initial text segment differs while the rest remains the same.
-Figure 2.b provides a schematic illustration of the output for the input shown in Figure 2.a.
-Because the UAT can only fit a single function, the output closely follows the trend of the input, resulting in nearly identical outcomes regardless of the slight differences in the initial part.
-Essentially, the output remains the same regardless of the prompt.
-In contrast, Figure 2.c shows the schematic output of the DUAT (Dynamic Universal Approximation Theorem) function.
-In DUAT, some or all parameters are functions of the input, allowing it to produce different outputs based on different prompts.
-This flexibility enables DUAT to generate varied results.
-On the right side, Figure 2 compares UAT and DUAT through the distribution of data.
-Since the parameters of UAT are fixed, it can only fit a fixed data distribution.
-Given that we typically train with large datasets, the goal for UAT is to fit as much data as possible, which often results in fitting only the densest part of the data distribution, represented in purple.
-On the other hand, DUAT can dynamically adjust its parameters based on the input, allowing it to fit multiple different distributions simultaneously.
-This comparison highlights the advantages of DUAT in handling more complex and diverse data scenarios, making it a more versatile and powerful tool for various applications.
+On the left side, Figure 2.a illustrates the limitation of the Universal Approximation Theorem (UAT) in fitting only a single function due to its fixed parameters post-training.
+In contrast, Figure 2.b demonstrates the capability of the Dynamic Universal Approximation Theorem (DUAT) to approximate multiple functions.
+In DUAT, some or all parameters are functions of the input, allowing it to dynamically adjust the parameters that would be fixed in UAT, thereby fitting different functions based on varying inputs.
+This adaptability enables DUAT to generate a diverse range of outputs.
+On the right side, Figure 2 compares the data distribution fitting capabilities of UAT and DUAT.
+Since UAT has fixed parameters, it can only fit a single, fixed data distribution.
+Given that training typically involves large datasets, the objective for UAT is to fit as much data as possible, which often results in fitting only the densest regions of the data distribution, highlighted in purple.
+Conversely, DUAT can dynamically adjust its parameters based on the input, allowing it to simultaneously fit multiple distinct distributions.
+This comparison underscores the advantages of DUAT in handling more complex and diverse data scenarios, making it a more versatile tool for various applications.
 3 DUAT for Transformer